### v1

# FakeScope: Large Multimodal Expert Model for Transparent AI-Generated Image Forensics

Abstract
The rapid and unrestrained advancement of generative artificial intelligence (AI) presents a double-edged sword: while enabling unprecedented creativity, it also facilitates the generation of highly convincing deceptive content, undermining societal trust. As image generation techniques become increasingly sophisticated, detecting synthetic images is no longer just a binary task—it necessitates interpretable, context-aware methodologies that enhance trustworthiness and transparency. However, existing detection models primarily focus on classification, offering limited explanatory insights into image authenticity. In this work, we propose FakeScope, an expert multimodal model (LMM) tailored for AI-generated image forensics, which not only identifies AI-synthetic images with high accuracy but also provides rich, interpretable, and query-driven forensic insights. To this end, we first construct FakeChain dataset that contains linguistic authenticity reasoning based on visual trace evidence, developed through a novel human-machine collaborative framework. Building upon it, we further present FakeInstruct, the largest multimodal instruction tuning dataset containing 2 million visual instructions tailored to enhance forensic awareness in LMMs. Leveraging the knowledge of FakeInstruct, FakeScope achieves state-of-the-art performance in both closed-ended and open-ended forensic scenarios. It can distinguish synthetic images with high accuracy while offering coherent and insightful explanations, free-form discussions on fine-grained forgery attributes, and actionable enhancement strategies. Notably, despite being trained exclusively on qualitative hard labels, FakeScope demonstrates remarkable zero-shot quantitative capability on detection, enabled by our proposed token-based probability estimation strategy. Furthermore, FakeScope exhibits strong generalization and in-the-wild ability, ensuring its applicability in real-world scenarios. The data, model, and demo will be publicly released on https://github.com/Yixuan423/FakeScope.
Index Terms:
AI-generated image detection, large multimodal models, forensic investigation, trustworthiness.
I Introduction
With the rapid development of image generation techniques, AI-synthesized images are becoming increasingly indistinguishable from genuine ones [1, 2]. While this prosperity has driven innovation, it also introduces significant cybersecurity threats, including misinformation, propaganda, and fraud. In response, substantial efforts have been made to detect AI-generated content in recent years [3, 4, 5, 6]. However, as generative models continue to evolve, concerns over the reliability and trustworthiness of detection systems in real-world applications persist. Addressing these challenges requires enhancing the transparency of AI-generated image detection models. We advocate for a paradigm shift in which image synthesis detection goes beyond binary classification, incorporating explainability to provide contextual insights into authenticity assisting trustworthiness.
Figure 1: Comparative illustration of image forensic model evolution. (a) Binary detectors; (b) general LMMs; and (c) FakeScope, which translates visual abnormalities into human-understandable cues—much like trace evidence in forensics—using natural language.
Figure 2: Contributions of this work. (a) FakeChain dataset, containing reasoning on image authenticity, constructed via the proposed ACoTI strategy (Sec. III); (b) FakeInstruct dataset, containing 2 million visual instructions of image forensic knowledge (Sec. IV); (c) FakeScope model, the expert model for transparent AI-generated image forensics, capable of multi-dimensional forensic capabilities (Sec. V).
Until recently, AI-generated image detection was predominantly framed as a binary classification task. Conventional approaches rely on unimodal architectures such as convolutional neural networks (CNNs) and vision transformers (ViTs) to classify images as either real or synthetic, significantly aiding in mitigating the spread of synthetic content. However, these early methods largely overlooked the need for transparency, offering little more than abstract heatmaps derived from deep neural network features as visual explanations. Recent research has shown that AI-generated images increasingly pass sanity checks by humans and machines [7, 1], leading to human-machine or machine-machine disagreements. As illustrated in Fig. 1 (a), different detectors can yield inconsistent decisions, significantly reducing practical utility. In contrast, in Fig.1(c), a transparent model can provide additional insights via visual trace evidence to enhance interpretability. Therefore, we anticipate more trustworthy paradigms that exceed binary detectors for achieving transparent forensics of AI-generated images. Such a model would enable human oversight, mitigate detection biases, and foster better confidence in AI-generated content forensics [2].
With the advent of multimodal models, visual-language approaches have started to emerge. Zhang et al. [8] introduced textual template descriptions to improve explainable deepfake face detection. Sun et al. [9] formulated brief descriptions by integrating forgery regions and candidate types using a predefined template. Li et al. [7] and Ye et al. [10] explored the capabilities of large multimodal models (LMMs) in the forensic analysis of AI-generated images. In [11, 12, 13], LMMs are incorporated to localize tampering forgeries and exhibits substantial effectiveness. Besides, LMMs are employed to assist in improving detection accuracy. In [14] and [15], conventional detectors are combined with LMMs to improve detection accuracy. Wen et al. [16] train a binary classifier with ViT intermediate features and use LMMs to explain the detection. These studies establish a concrete foundation for the transition from unimodal to multimodal approaches in the area of AI-generated image forensics. Nevertheless, due to the inherent limitations of LMMs as classifiers [17], previous approaches have struggled to balance detection accuracy and transparency, often relying on separate strategies and training procedures to fulfill multi-dimensional forensic tasks. Therefore, there is an urgent need for a unified forensic expert model for AI-generated images capable of multiple forensic tasks in diversified scenarios.
Inspired by the above insights, we present FakeScope, a unified expert model for transparent and versatile AI-generated image forensics. It leverages the remarkable cross-modality ability of generally pretrained LMMs while being systematically enhanced for universal forensic awareness. Specifically, it is designed with three key objectives: (1) comprehension on trace evidence and authenticity; (2) faithful response to diverse user queries; and (3) qualitative and quantitative detection without model extension. To achieve these, we collect and leverage forensic knowledge to improve LMMs’ forensic awareness on AI-generated images. As shown in Fig. 2(a), we first construct FakeChain, a large-scale multimodal dataset containing long and comprehensive reasoning on image authenticity as forensic knowledge source, developed via a cost-efficient human-in-the-loop strategy. We further enrich FakeChain to create FakeInstruct, a gigantic-scale dataset comprising 2 million visual instructions to enhance diverse forensic abilities of LMMs (Fig. 2(b)). Extensive experiments demonstrate that FakeInstruct consistently improves the forensic capabilities of general LMMs, and the proposed FakeScope (Fig. 2(c)) exhibits state-of-the-art (SOTA) performance in detecting, explaining, analyzing, and discussing AI-generated images. In particular, it can generalize well to unseen image content, generation models, and even user-created contents. Our core contributions are as follows.
  • We present FakeChain, a large-scale multimodal dataset with coherent reasoning flow from visual trace evidence to image authenticity. It is collected through our proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) scheme, a cost-effective human-machine collaboration strategy that requires minimal human annotation while ensuring reliability.
  • We develop FakeInstruct, the first million-scale dataset designed to enhance LMMs’ multi-dimensional forensic capabilities in AI-generated image detection.
  • We propose FakeScope, a multimodal expert model for transparent AI-generated image detection that integrates both detection and query-based forensic analysis. We further propose a zero-shot token-based strategy for estimating detection probabilities. Extensive experiments validate FakeScope’s superiority in both detection and transparency-related tasks, outperforming other LMMs and binary models across diverse datasets.
II Related Works
Recently, the rapid advancement of image generation technology has driven growing interest in AI-generated image forensics, and the evolution of LMMs has opened new avenues for enhancing the trustworthiness of image forensic systems. We first outline the progression of AI-generated image detection techniques, followed by an overview of the current developments in LMMs.
II-A AI-generated Image Detection
The detection of AI-generated images aims to differentiate fake images (synthesized by generative AI algorithms) from real ones (created or captured by humans). This problem is commonly formulated as a binary classification task. Preexistent approaches generally incorporate forensic patterns to classify, which can be categorized into the flaw-based [18, 19] and fingerprints-based [20, 21, 22, 23] according to their forensic ground. The flaw-based methods capture the intricate details in low (e.g., texture [18]) or high (e.g., physical rules [19, 24]) semantic levels, while the fingerprints-based detectors identify the invisible leftover patterns of generators (e.g., frequency artifacts [20, 21, 22, 23]). Moreover, investigations on model’s generalization ability towards unseen  [25, 26, 4, 3, 5, 6] and generator attribution [27, 28] have emerged. Some of them assume to be only accessible to real images and utilize generators (like an autoencoder) to simulate generative artifacts [29, 22, 26, 25]. Although remarkable, current fake image detection models still fall short in the interpretability towards human [30], which causes the detection results hard-to-explain for the general populace and weakens the credibility of model output. Therefore, developing an expert dedicated to fake image detection with both high detection accuracy and transparency is currently urgently needed.
II-B Large Multimodal Models
Typical LMMs compromise a modality encoder, a large language model (LLM), with a modality interface for cross-modal connection [31, 32]. The specialty of LMM lies in their ability to effectively receive, process, and generate multimodal information. Currently, LMMs are developing by leaps and bounds and have exhibited remarkable visual-language capabilities on par with or even surpassing the human, including the closed-source models (e.g., GPT-4V [33], Gemini [34], and Claude [35]) and the open-source ones (e.g., LLaVA [36], mPLUG-Owl [37], deepseek [38]). Furthermore, to adapt general-purpose LMMs to specialized fields, recent researchers proposed various expert models based on visual instruction tuning  [39, 40, 41, 36]. Despite great progress, recent studies [7] indicate that LMMs’ forensic capabilities are still unsatisfactory, especially when it comes to identifying fine-grained forgery aspects. On the other hand, human crowd-sourcing data for visual instruction tuning becomes increasingly hard to get, partially due to the process being time-consuming, especially when multimodal information is considered. In this work, to infuse the forensic abilities on AI-generated images into foundational LMMs, we construct the first visual instruction tuning dataset centering on fake image detection, the FakeInstruct, aiming at a unified multimodal paradigm based on LMMs for more transparent image forensics.
III FakeChain: Reasoning on Image Authenticity
In this section, we elaborate on the construction of FakeChain, a large-scale dataset with long and comprehensive reasoning on underlying trace evidence for authenticity. It functions as the source of forensic knowledge and further contributes to building the instruction dataset and the forensic model. For construction, we first prepare images (Sec. III-A) and then collect authenticity reasoning. While human annotation remains the most intuitive and widely applied approach for training LMMs [42, 36], acquiring sufficiently fine-grained dataset for authenticity reasoning is highly labor-intensive. To mitigate the prohibitive costs, a natural idea is to use a human-machine collaboration strategy. Therefore, we first collect human expert feedback (Sec. III-B) as weak supervision in the proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy (Sec. III-C) for forensic reasoning collection.
Figure 3: Pipeline of the proposed ACoTI strategy for obtaining the reasoning on trace evidence under the human-machine collaboration scheme. ACoTI is cost-efficient that adopts the human-in-the-loop methodology to incorporate little human labor to demonstrate, steer, and enlighten non-perfect AI experts to generate qualified reasoning data. This semi-automatic strategy unburdens human labor, leverages LMM’s abundant inner knowledge, and ensures reliability simultaneously.
III-A Image Preparation
FakeChain contains both AI-generated and real images in equal quantities. In contrast to the one-to-all generalization setting in some works [25, 26, 4, 3, 5, 6], FakeChain is deliberately constructed with a high degree of diversity to capture a broad range of generative patterns. In particular, AI-generated images function as the knowledge source of forensic trace evidence. To ensure content diversity, we sampled 23,797 AI-generated images from multiple sources. These generators span 17 types, including GAN-based, diffusion-based, auto-regressive, and proprietary models. In specific, we choose representative ProGAN [43], StyleGAN [44], BigGAN [45], CogView2 [46], ADM [47], IF [48], FuseDream [49], VQDM [50], Glide [51], SD [52], SDXL [53], FLUX.1 [54], Firefly [55], Dalle2 [56], Dalle3 [57], Wukong [58] and Midjourney [59]. Also, we collect 23,797 real images from DIV2K [60], RAISE [61], and ImageNet [62]. The details of image components are provided in the supplementary material (SM).
III-B Human Expert Feedback
As shown in Fig. 3, we require a few human annotations in step1-steer and step2-demonstrate of ACoTI, including the trace evidence category and reasoning demonstration respectively. Therefore, we conduct a well-controlled in-lab experiment involving 30 trained human experts with experience in image generation as annotators, ensuring qualified understanding of generative patterns. For step1-steer, experts are required to correctly distinguish whether the image is AI-generated, and then select supporting categories, ensuring only accurate judgments contribute to the dataset. In practice, we provide a category set for reference $\mathcal{F}$ = {texture, edge, clarity, light&shadow, anatomy, layout, symmetry, reflection, perspective, physics, shape, theme, content deficiency, distortion, unrealistic, overall hue[7], which can characterize authenticity across different levels of trace evidence. Each image is annotated by two and only the intersection of selected categories is retained to improve reliability. Additionally, experts can provide supplementary authenticity-related information at will, such as specific area of occurrence and commonsense knowledge. For step 2-demonstrate, we fix 50 fake and 50 genuine images, and employ experts to annotate manually with detailed reasoning. These expert-annotated images constitute the exemplar sets $\{{\chi}^{+},{\chi}^{-}\}$ as illustrated in upper Fig. 3, where + and - denote AI-generated and genuine images respectively. To ensure balance, we manually guarantee each category of $\mathcal{F}$ emerges in at least three exemplars. Herein, we establish the necessary human supervision for the subsequent reasoning collection.
III-C Reasoning on Visual Trace Evidence
III-C1 Task Definition
FakeChain encapsulates forensic knowledge to enable models a fundamental understanding between visual trace evidence and authenticity. We structure it as a materialized reasoning process, comprising an exhaustive description of the appearance and location of trace evidence, enclosed with an authenticity conclusion. Compared with free-form descriptions, this structured reasoning format has two advantages: (1) it allows for a precise and detailed reflection of the aspects humans focus on when assessing authenticity; and (2) it inherently follows the logical pattern of human thought. For instance, when examining a suspicious image, humans first scrutinize telltale visual clues and then summarize an inclined conclusion like “Thus, I tend to think this is a fake image generated by AI”. This structured reasoning approach helps LMMs more effectively mimic human decision-making in AI-generated image detection, thereby enhancing human-centered transparency.
Figure 4: The composition of the FakeInstruct which is derived from the FakeChain dataset, containing 47K visual instructions on absolute authenticity judgments, 95K instructions on holistic reasoning, 715K on fine-grained visual trace evidence, and 1190K on extended knowledge and discussions. The million-scale diversified visual instructions of FakeInstruct enable LMMs with both broad and fine-grained understanding of image authenticity, ensuring LMMs are equipped to handle diverse forensic tasks with both interpretability and accuracy.
III-C2 ACoTI Strategy
To develop a cost-efficient approach for obtaining large-scale reasoning data, we present a semi-automatic scheme to efficiently extract forensic knowledge from AI experts with few human annotations as supervision, i.e., the Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy. As illustrated in Fig. 3, ACoTI operates in a human-machine collaborative framework, where humans partially participate in the first and second phases by providing partial information of trace evidence. Inspired by human reasoning processes, ACoTI follows the Steer-Demonstrate-Enlighten pipeline, with the entire process detailed as follows.
Step 1. Steer: To utilize off-the-shelf LMMs maximally on reasoning, we employ the category of visual trace evidence to steer the reasoning process. We first notify the AI expert with the human-annotated category set $\mathcal{F}_{i}$ for the $i^{th}$ target image to ensure reasoning process focused on forensic trace evidence rather than drifting into unrelated domains. The $\mathcal{F}_{i}$ directly participates in the following two steps.
Step 2. Demonstrate: To further mitigate hallucinations in imperfect models, we leverage human demonstrations as in-context examples, inspired by LMMs’ ability to learn from explanations in context [63]. Considering the trade-off between context length and efficiency, we adaptively select the most relevant demonstrations by filtering a minimum covering set [64] ${\chi}_{i}$ from $\{{\chi}^{+},{\chi}^{-}\}$ for $\mathcal{F}_{i}$. We ensure at least one positive sample and one negative sample are included in ${\chi}_{i}$ via optimizing the following process:
$\chi_{i}=\arg\min_{\chi^{\prime}\subseteq\{\chi^{+}\cup\chi^{-}\}}|\chi^{\prime}|,$ (1)
[…]
where $L(\cdot)$ represents the mapping function that extracts the categories covered by a given subset of exemplars. This ensures the selected examples can comprehensively represent the key forensic attributes required for reasoning.
Step 3. Enlighten: The core of ACoTI lies in enlightening the AI expert through a carefully structured prompting scheme. As illustrated in Fig. 3, the proposed ACoTI prompting emulates human reasoning process with few-shot chain-of-thought (CoT) prompting [65]. A unique aspect is explicitly providing the AI with the correct authenticity label in the context, guiding the model toward the appropriate answer space [65]. Finally, this framework is structured as a quadruple: <context, target, human supervisions, CoT prompt>, where human supervision contains $\mathcal{F}_{i}$, $\chi_{i}$ and $auth_{i}$. Detailed prompts are provided in the SM.
In this way, ACoTI provides an effective pipeline for distilling forensic knowledge from strong models into usable reasoning data. Compared with manually collecting long and complex descriptions from human annotators, ACoTI significantly reduces the burden while maintaining data quality. To quantify its effectiveness, we further conducted a two-alternative forced choice (2AFC) study on human preference between the outcomes generated with ACoTI and those without. Six annotators rated on 100 randomly sampled annotated images, and ACoTI demonstrated 99.17% preference, indicating its effectiveness as a cost-efficient approach for extracting forensic knowledge from imperfect models. Therefore, we apply ACoTI to all collected images and organize the resulting data into the predefined reasoning format $rsn$. As shown in Fig. 4(b), we provide an image with corresponding $rsn$. The 47,594 tuples of $\{\bm{\mathtt{I}},auth,rsn\}$ compose the FakeChain dataset, where $\bm{\mathtt{I}}$ and $auth$ denote the image and authenticity label respectively.
IV FakeInstruct: Million-scale Multimodal Forensic Instruction Dataset
Although FakeChain offers corpus-rich reasoning with abundant forensic knowledge, we transform it into a diverse set of visual instructions to better leverage the information embedded in long-form text. As illustrated in Fig. 2(b), we further derive instruction-following data from FakeChain in the visual question answering (VQA) format with both Mistral [66] and GPT-4o [33] in an intermingling manner to enhance linguistic and instructional diversity. We take extra precautions to remove textual contamination allowing models to infer answers without visual information, thereby ensuring the integrity of instruction-based reasoning [67]. As depicted in Fig. 4, the dataset contains 2000K diversified instructions in total, which endows LMMs with flexibility and knowledge in handling extensive forensic scenarios. The details are elaborated as follows.
Part 1. Absolute Authenticity Judgments: As shown in Fig. 4(a), we design diverse question formulations on absolute image authenticity to explicitly implant the concepts of fake and real in LMMs. Throughout this part, we utilize the term fake to denote the AI-generated images, while real for the opposite. While incorporating synonyms such as genuine, fault, and artificial can enhance interaction flexibility with human users, we prioritize a higher level of certainty in objective authenticity assessment in this subset.
Part 2. Holistic Reasoning: As shown in Fig. 4(b), the holistic reasoning imitates human inference on image authenticity, which can foster LMMs’ awareness of the relationships between multi-level visual trace evidence and the overall authenticity. Additionally, the carefully structured cause-and-effect reasoning framework enhances LMMs’ ability to distinguish AI-generated images by activating their latent reasoning capabilities [68].
Part 3. Fine-grained Trace Evidence: To equip LMMs with the ability to seamlessly respond to diverse forensic queries from users, we decompose and reformat the reasoning data in FakeChain into a variety of visual instructions, structured in narrative or multiple-choice question (MCQ) [69] formats. This transformation is facilitated by GPT-4o [33] and Mistral [66], whose complementary strengths ensure a diverse distribution of instructional corpus. As illustrated in Fig. 4(c), the dataset includes authenticity-related VQAs covering what, how, yes-or-no, and imperative question types. These questions pertain to perceptual adjectives (e.g., unnatural, misaligned), questionable visual attributes (e.g., smudge, shape), regions of occurrence (e.g., right eye, apple skin), or contextual knowledge (e.g., shadow in wrong directions). The MCQs comprise one correct answer alongside multiple plausible yet misleading distractors, thereby enhancing models’ ability to discern subtle visual cues while mitigating misconceptions.
Part 4. Extensional Knowledge: Beyond fundamental trace evidence reasoning and forensic analysis, we further incorporate extensional knowledge through open-ended discussions (Fig. 4(d)). Specifically, these discussions cover four major aspects: origins of perceptual misjudgments arising from trace evidence, suggesting improvements to enhance the realism of generation, inferring potential generator characteristics based on forensic evidence, and exploring additional user-driven inquiries related to authenticity.
We herein present a brief analysis of the FakeInstruct dataset. In total, about 2000K visual instructions are included in the FakeInstruct dataset. As shown in Fig. 4(e), quantitatively, the image content type and instruction type are well-balanced. Qualitatively, Fig. 4(c) shows decent diversity and information consistency across instructions.
V FakeScope: Large Multimodal Forensic Model
In this section, we introduce the proposed large multimodal expert model, FakeScope, which leverages state-of-the-art LMMs as its backbone and is further fine-tuned on FakeInstruct to incorporate forensic knowledge. Moreover, we propose a non-trivial probability estimation strategy via token soft-scoring, which could extend the model’s qualitative authenticity judgments to quantifiable probability estimations, without requiring additional explicit training on numerical labels. The details are elaborated as follows.
V-A Architecture and Supervised Instruction Tuning
The architecture of FakeScope follows the LLaVA-style [36] vision-language model (VLM) framework, compromising an image encoder that transforms images into tokenized representations, an LLM for textual interpretation and generation, and a vision-language connector that bridge modalities. As listed in Table I, to validate the universality of FakeInstruct, we incorporate two different model architectures, including the 7B scale LLaVA-v1.5 [36] and mPLUG-Owl2 [37], corresponding to two variants of FakeScope. LMM training typically involves two phases: the cross-modal alignment requiring gigantic-scale data and visual instruction tuning on data of proprietary domain [36]. Given the extensive scale of FakeInstruct, we introduce an auxiliary training stage to perform supervised visual instruction tuning specifically on FakeInstruct. Verified by previous studies [70], adequate tuning on the visual instructions in FakeInstruct is expected to facilitate the model’s visual forensic capabilities while also preserving the general knowledge obtained from earlier-phase training. To optimize efficiency, we freeze the vision encoder and fine-tune only the modality projector and language model.
V-B Probability Estimation via Token Soft-Scoring
Figure 5: Probability estimation in FakeScope. The log probabilities of candidate tokens at the <auth_token>  position are extracted and normalized using softmax weighting to compute the predicted authenticity probability. The illustration is based on the mPLUG-Owl2 [37] backbone.
TABLE I: Meta structures of LMMs as FakeScope foundation.
[…]
LLaVA-v1.5 (7B) [36] CLIP-ViT-L14 MLP Vicuna-7B
mPLUG-Owl2 [37] CLIP-ViT-L14 Abstractor LLaMA2-7B
TABLE II: Publicly available datasets used for evaluation. The ‡ notion indicates the uncontaminated content of the dataset, and # denotes the unit number.
Purpose Datasets Modality # Real/ # Fake Real Image Source # Generators Evaluation Criteria
Multimodal Forensic Capability FakeBench [7] Multiple 3k / 3k ImageNet, DIV2K, COCO 10 Detecting, Interpreting, Reasoning, Fine-grained Analyzing
Generalization Capability AntiFakePrompt [71] Single 42k / 6k COCO 11 Detecting
AIGCDetectBenchmark [72] 74.3k / 74.3k COCO, LSUN, FFHQ, CelebA 7
In-the-wild Performance WildRF [73] Single 1.25k / 1.25k Reddit, FB, X Unknown Detecting
SynthWildX [4] 500 / 1.5k X 3
LMMs typically convey their forensic judgment through hard token selection, i.e., choosing the token with the highest logit when generating responses. Here, “token” refers to the vocabulary-level prediction unit in the model’s output distribution, with each token associated with a logit indicating its selection likelihood. This approach inherently limits their ability to provide numerical probability scores, which are crucial for assessing model confidence. Moreover, general-purpose LMMs are not explicitly trained as classifiers with numerical labels, making direct probability estimation an unpreferred methodology. To address this limitation, FakeScope is trained using binary authenticity judgments and incorporates a token soft-scoring strategy to convert model outputs into quantitative probability estimates. In specific, as depicted in Fig. 5, we define fake and real as anchor tokens and apply softmax normalization over their corresponding logits within a standardized prompting framework:
USER: <Image> What is the authenticity of this image?
ASSISTANT: This image is [auth_token].
To compute probabilities, we extract the logits at the specific position of [auth_token], denoted by $\mathcal{T}=\{\tilde{t}_{j}|_{j=0}^{1}\}=\{\tilde{t}_{fake},\ \tilde{t}_{real}\}$, where $\tilde{t}_{(\cdot)}$ represents case variations of the anchor tokens to enhance robustness. The probability $p_{i}$ of each class is then calculated as follows,
$p_{j}=\frac{e^{\sum_{v\in\tilde{t}_{j}}z^{(v)}}}{\sum_{j=0}^{1}e^{\sum_{v\in\tilde{t}_{j}}z^{(v)}}},$ (2)
where $z^{(v)}$ represents the logit associated with token $v$. By leveraging token-level probability distributions rather than solely selecting the most probable token, this approach retains nuanced confidence signals that would otherwise be discarded during deterministic text generation. This enables FakeScope to generate quantifiable probability estimates, improving its interpretability as a forensic detector.
VI Experiments and Analyses
VI-A Protocols
In this section, we present a systematic evaluation of FakeScope, analyzing several key aspects of its performance:
  • Multimodal Forensic Capabilities: We evaluate to what extent the FakeScope can function as a transparent image forensic model (Sec. VI-B).
  • In-the-wild Performance: We assess FakeScope’s performance on user-generated images collected from social networks (Sec. VI-C).
  • Generalization Capability: We examine how well FakeScope generalizes to unseen image content and generators (Sec. VI-D).
Additionally, we perform ablation studies to analyze the impact of different training data components and scales (Sec. VI-E). Detailed experimental setups are provided below.
VI-A1 Implementation Details
Training FakeScope is implemented with PyTorch on eight Nvidia H20-NVLink 98G GPUs, and the evaluations are conducted on one NVIDIA Tesla A100 40G GPU. All training are set to full-schedule mode, with detailed parameter configurations provided in SM. To ensure FakeScope serves as a unified forensic model for AI-generated images, we retain a single checkpoint per base model for all evaluations, ensuring fair comparisons. Furthermore, to prevent data leakage, no images from the testing datasets are reused in the training data. To ensure fairness, no data augmentation strategy is employed throughout the entire training process of FakeScope.
VI-A2 Evaluation Datasets
We evaluate the proposed FakeScope’s multifaceted performance on five commonly used datasets in the area of AI-generated image detection. As listed in Table II, we validate multimodal forensic capabilities on FakeBench [7], covering both binary authenticity and multimodal forensic criteria. To assess generalization to unseen content and generation models, we use AntifakePrompt [71], which contains fake images derived from real Microsoft COCO [74] images that are unseen during training. Additionally, we employ the non-overlapping subset of AIGCDetectBenchmark [72], excluding samples overlapping with ImageNet [62] to avoid contamination and ensure fair generalization evaluation. We further test FakeScope on user-generated images from social networks to support in-the-wild evaluation, using the WildRF [73] and SynthWildX [4] datasets, where authenticity labels are derived from social tags. It is worth noting that the scenarios like deepfakes (e.g. the FF++ dataset [75]) and are treated as out-of-distribution (OOD) cases, given their fundamental difference from fully AI-generated images. Nonetheless, they are included in our generalization evaluation to assess the model’s adaptability.
VI-A3 Baseline Models
The proposed FakeScope is compared with a diverse set of models, including open-sourced LMMs (InstructBLIP [76], IDEFICS-Instruct [77], InternLM-XComposer2-vl [31], LLaVA-v1.5 [36], Qwen-VL [78], mPLUG-Owl2 [37]), proprietary LMMs (GPT-4V [33], GeminiPro [34], Claude3-Sonnet [35]), data-driven binary forensic models (CNNSpot [20], UnivFD [3], FreDect [22]), and the vision-language contrastive model CLIP-ViT-Large-14 [79]. For detection evaluation, LMMs supports both qualitative (textual judgments) and quantitative (probabilistic estimation) settings, whereas binary detectors are limited to the quantitative mode. To ensure fair comparisons, the data-driven models are retrained on the same images as FakeScope without data augmentation, and all tested LMMs are evaluated using same prompting templates as FakeScope. These settings remain consistent across all experiments unless explicitly stated otherwise.
TABLE III: Forensic detection results on the FakeBench-FakeClass [7] dataset under the qualitative setting. The detection Acc (%) of each model is reported. The first and second best LMMs are highlighted in bold and underlined, respectively. The “Fake” column reports the average accuracy across all individual “Generation Model” columns, and both the averaged “Authenticity” columns and “Question Type” columns reflect the “Overall” column.
Subcategories Authenticity Question Type Generation Model Overall
[…]
Random guess 42.27 58.03 51.80 48.50 40.33 36.00 51.67 33.67 38.00 44.67 32.33 47.67 37.67 60.67 50.15
Human (Best) 97.00 93.00 / / 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.00 87.50
Human (Worst) 43.00 13.00 / / 50.00 0.00 46.15 77.78 33.33 16.67 41.67 7.14 14.29 0.00 55.00
Human (Overall) 76.91 72.12 / / 95.47 58.82 87.07 95.90 63.96 82.91 79.41 77.66 70.69 45.59 74.51
GPT-4V (Proprietary, teacher) 59.87 96.20 76.33 79.73 95.00 66.00 66.67 81.33 36.67 49.00 45.67 44.33 61.00 53.00 78.03
GeminiPro (Proprietary) 35.83 99.27 61.27 73.83 45.33 15.67 47.33 52.67 19.33 28.00 35.33 35.67 41.00 38.00 67.50
Claude3 Sonnet (Proprietary) 12.00 98.23 55.97 54.27 7.00 0.00 20.33 13.67 0.67 2.67 15.33 9.33 30.33 20.67 55.12
InternLM-XC.2-vl (InternLM2-7B) 32.17 92.33 60.40 64.10 54.00 12.67 43.33 49.00 19.00 20.33 29.33 29.67 34.67 29.67 62.25
InstructBLIP (Vicuna-7B) 67.80 47.67 65.23 50.23 80.67 65.33 66.00 71.33 68.00 70.33 60.33 61.00 70.67 64.33 57.73
Qwen-VL (Qwen-7B) 28.57 84.27 51.17 61.17 45.67 5.67 35.00 37.67 8.33 14.33 31.33 21.00 46.67 40.00 56.42
IDEFICS-Instruct (LLaMA-7B) 24.97 31.97 7.00 49.93 19.67 27.33 24.33 28.33 27.67 24.00 23.67 23.67 24.00 27.00 28.47
LLaVA-v1.5 (Vicuna-7B, baseline) 38.00 77.40 55.77 59.63 44.33 17.33 45.33 55.33 23.00 28.33 39.67 0.00 0.00 41.00 57.70
mPLUG-Owl2 (LaMA2-7B, baseline) 49.60 93.97 71.87 71.70 56.33 10.00 72.00 75.33 19.33 32.00 56.00 55.00 64.00 56.00 71.78
FakeScope (Ours, mPLUG-Owl2) 99.50 99.87 99.73 99.63 100.00 100.00 99.67 100.00 100.00 100.00 98.67 100.00 97.67 99.00 99.68
Improvement +49.90 +5.90 +27.86 +27.93 +43.67 +90.00 +27.67 +24.67 +80.67 +68.00 +42.67 +45.00 +33.67 +43.00 +27.90
FakeScope (Ours, LLaVA-v1.5) 98.67 97.93 98.67 97.93 99.67 98.67 99.00 99.67 96.67 97.00 99.33 100.00 98.00 98.67 98.30
Improvement +60.67 +20.53 +42.90 +38.30 +55.34 +81.34 +53.67 +44.34 +73.67 +68.67 +59.66 +100.00 +98.00 +57.67 +40.60
VI-A4 Performance Measures
FakeScope is versatile in coping with forensic detection and in-depth analysis simultaneously. To comprehensively evaluate its multidimensional capabilities, we adopt a range of metrics. For forensic detection, we consider both qualitative and quantitative settings. Specifically, qualitative classification accuracy is measured by $Acc=N_{c}/N_{t}$, where $N_{c}$ and $N_{t}$ denote the number of correct and total predictions, respectively. For quantitative assessment, we follow [20, 3] and use threshold-independent calibration metrics, including average precision (AP) and area under the receiver operating characteristic curve (AUC), which are robust to class imbalance across datasets. To assess transparency-related performance, we evaluate the alignment between model responses and reference answers by employing both conventional word-level metrics (i.e., BLEU [80], ROUGE [81], and BERTScore [82]) and the LLM-as-a-judge strategy [83] to measure semantic-level alignment, following [40, 36, 7, 16]. The measures are then normalized and averaged to obtain the overall reflection, where higher values indicate better performance on transparency-related abilities. Details on evaluation measures are provided in SM.
Figure 6: Forensic detection results on FakeClass [7] under the quantitative setting. Calibration measures Acc (threshold is set to 0.5 for balanced indication), AP, and AUC (%) are reported for each model. The FakeScope far exceeds the other models on each subset and the entire dataset.
VI-B Multimodal Forensic Capability
The multimodal forensic capabilities of the proposed FakeScope model are quantitatively assessed on FakeBench containing FakeClass, FakeClue, and FakeQA subsets with respect to four ability criteria for LMMs comprehensively defined in [7], which are stated as follows.
  • Detecting, which evaluates whether forensic classifiers can accurately differentiate AI-generated fake images from real ones;
  • Interpreting, which examines how well LMMs can extract visual evidence to support the known authenticity;
  • Reasoning, which assesses whether LMMs can identify generative visual trace evidence and correctly infer the unknown image authenticity;
  • Fine-grained Analyzing, which investigates whether LMMs can discuss and analyze specific aspects of telltale visual clues.
In practice, all LMM responses are generated using the greedy search strategy to ensure reproducibility. The quantitative evaluation results for the four criteria are presented below, with qualitative examples provided in SM.
TABLE IV: Multimodal forensic capability comparisons on FakeBench [7], including detecting on FakeClass, interpreting on FakeClue-faultfinding, reasoning on FakeClue-inference, and fine-grained forgery analyzing on FakeQA. The B.-1/B.-2 denote BLEU-1/2, R.-L is ROUGE-L, Sim. is BERTScore similarity, and $\alpha$, $\rho$, $\kappa$ correspond to completeness, preciseness, and relevance judged by LLMs. The first- and second-best performances of each column are highlighted in bold and underlined, respectively.
Capabilities Interpreting Reasoning Fine-grained Analyzing Overall
LMMs Word-level Semantic-level Avr.$\uparrow$ Word-level Semantic-level Avr.$\uparrow$ Word-level Semantic-level Avr.$\uparrow$
B.-1 B.-2 R.-L Sim. $\alpha$ $\rho$ $\kappa$ B.-1 B.-2 R.-L Sim. $\alpha$ $\rho$ $\kappa$ B.-1 B.-2 R.-L Sim. $\alpha$ $\rho$ $\kappa$
GPT-4V (Teacher) 0.188 0.092 0.211 0.635 1.872 1.423 1.891 0.573 0.273 0.118 0.225 0.612 1.392 0.777 1.503 0.460 0.062 0.022 0.177 0.455 1.518 1.261 1.869 0.477 0.503
GeminiPro 0.068 0.029 0.113 0.459 1.688 1.024 1.271 0.415 0.168 0.059 0.181 0.485 0.833 0.561 1.239 0.331 0.178 0.071 0.199 0.365 1.247 0.675 1.686 0.402 0.383
Claude3 Sonnet 0.130 0.042 0.170 0.467 1.516 1.182 1.884 0.483 0.188 0.062 0.193 0.528 0.928 0.441 1.456 0.357 0.038 0.012 0.158 0.398 1.254 1.091 1.886 0.429 0.423
InstructBLIP 0.207 0.072 0.206 0.463 1.841 1.398 1.829 0.541 0.166 0.065 0.160 0.458 1.828 0.922 1.546 0.464 0.004 0.001 0.007 0.079 0.810 0.321 0.543 0.151 0.385
IDEFICS-Instruct 0.093 0.032 0.140 0.367 1.936 1.344 1.789 0.501 0.099 0.026 0.148 0.457 1.178 0.450 0.967 0.308 0.189 0.086 0.231 0.381 1.165 0.456 1.483 0.370 0.393
InternLM-XC.2-vl 0.095 0.032 0.142 0.402 1.899 1.260 1.534 0.475 0.108 0.032 0.169 0.492 1.004 0.670 1.167 0.337 0.146 0.054 0.173 0.366 1.208 0.719 1.623 0.388 0.400
LLaVA-v1.5 0.143 0.047 0.159 0.466 1.812 0.964 1.621 0.468 0.151 0.051 0.167 0.465 0.905 0.404 1.166 0.311 0.128 0.051 0.161 0.360 1.259 0.757 1.659 0.394 0.391
Qwen-VL 0.056 0.018 0.096 0.322 1.853 1.149 1.485 0.432 0.130 0.045 0.159 0.469 0.929 0.401 1.104 0.303 0.076 0.026 0.105 0.241 1.004 0.490 0.802 0.247 0.327
mPLUG-Owl2 0.099 0.032 0.146 0.455 1.721 0.839 1.415 0.423 0.109 0.034 0.160 0.458 0.843 0.500 1.070 0.296 0.187 0.091 0.228 0.439 1.089 0.412 1.504 0.368 0.362
FakeScope-M 0.434 0.256 0.371 0.704 1.897 1.608 1.901 0.638 0.439 0.264 0.383 0.747 1.886 1.265 1.775 0.614 0.290 0.174 0.293 0.603 1.628 1.356 1.857 0.540 0.597
Improvement +0.335 +0.224 +0.225 +0.249 +0.176 +0.769 +0.486 +0.215 +0.330 +0.230 +0.223 +0.289 +1.043 +0.765 +0.705 +0.318 +0.103 +0.083 +0.065 +0.164 +0.539 +0.944 +0.353 +0.172 +0.235
FakeScope-L 0.432 0.254 0.370 0.708 1.893 1.653 1.887 0.640 0.437 0.262 0.381 0.745 1.873 1.242 1.776 0.610 0.304 0.184 0.305 0.611 1.647 1.402 1.879 0.553 0.601
Improvement +0.289 +0.207 +0.211 +0.242 +0.081 +0.689 +0.266 +0.172 +0.286 +0.211 +0.214 +0.280 +0.968 +0.838 +0.610 +0.299 +0.176 +0.133 +0.144 +0.251 +0.388 +0.645 +0.220 +0.159 +0.210
1) Detection Task: Simple detection of fake and real images is the fundamental capability of forensic detectors. In Table III, we compare the performance of the proposed FakeScope with other LMMs under the qualitative setting on the FakeClass dataset. The FakeClass dataset contains 6,000 diversified query-answer pairs for judging absolute image authenticity with balanced authenticity proportions and adequate query diversity. The detection accuracy is measured by the correctness of responses to these queries. As reflected in the results, FakeScope of the two versions simultaneously exhibits noticeable improvements in detection reliability across all generators, authenticity categories, and query types, suggesting that FakeInstruct can significantly enhance the forensic awareness of LMMs on binary classification. Notably, FakeScope achieves a peak accuracy of 99.68%, outperforming even the best-performing human expert, and marking a significant milestone in automated forensic analysis. In Fig. 6, we further present the performance comparison using calibration measures under the quantitative setting. In this case, one can still witness a consistently significant advantage across all categories over both other LMMs, indicating the effectiveness of the proposed probability estimation strategy based on token soft-scoring. Notably, since we do not explicitly incorporate numerical values as the supervision when training LMMs as forensic detectors, the results further suggest that LMMs can effectively adapt to hard authenticity labels via the internalized probabilistic reasoning process and perform well on estimating numerical probabilities. In particular, despite being trained with the same image data, CNNSpot [20], UnivFD [3], and FreDect [22] exhibit inferior performance compared with FakeScope. While these data-driven methods can be effective in some scenarios, their performance often benefits from elaborate data augmentation [20]. In contrast, under the controlled setting of this study where no image-level augmentation is applied, purely training with binary labels may fall short in capturing the intricacies of forensic cues. This further underscores the effectiveness of the multimodal information in FakeInstruct and highlights LMMs’ superiority to integrate multimodal information for AI-generated image forensics.
2) Interpreting Task: We find out that incorporating visual trace evidence into training can significantly enhance the interpretation ability of LMMs. In practice, LMMs are evaluated using the fault-finding prompting mode of FakeClue, where models are instructed to explain image authenticity based on visual clues, such as in the prompt “This is a fake image generated by AI, explain the reasons”. In the first category of Table IV, the two variants of FakeScope demonstrate substantial improvements in interpreting ability, especially the word-level similarities and semantic relevance level to the ground truth. These findings indicate that training on FakeInstruct conspicuously enhances the effect-to-cause awareness concerning the forensic attributes of LMMs.
3) Reasoning Task: Forensic reasoning is the inverse of interpretation, assessing LMMs’ inference capability in deriving authenticity judgments from visual trace evidence. This task is evaluated using the prompts in the inference mode of FakeClue, such as “Inspect this image carefully and formulate an evaluation regarding its authenticity”. As observed in the experimental results listed in the second category of Table IV, while all baseline LMMs generally exhibit weak forensic reasoning capabilities, the proposed FakeScope achieves a substantial improvement after training on FakeInstruct and even surpasses all compared models, reaching SOTA performance. Most notably, its reasoning capability also outperforms the teacher model GPT-4V [33] by over 30%, demonstrating the effectiveness of FakeInstruct in enhancing cause-to-effect forensic awareness and triggering forensic reasoning in LMMs.
4) Fine-Grained Analyzing Task: To further evaluate the effectiveness of FakeScope in analyzing specific aspects of image forgery, we conduct experiments on the FakeQA dataset, which consists of queries regarding fine-grained visual forgeries. As shown in the third category of Table IV, FakeScope exhibits notable improvements over both baseline models and its teacher model, highlighting its enhanced fine-grained forensic analyzing capabilities. However, despite these promising results, this ability still shows room for improvement compared to the other two transparency-related forensic capabilities interpreting and reasoning. This suggests potential avenues for further optimization, particularly in refining its sensitivity to subtle forgery artifacts.
TABLE V: In-the-wild forensic detection ability on WildRF [73] and SynthWildX [4] containing user-generated image content and unknown generators. The AP and AUC (%) are reported, and the first- and second-best performers for each category are highlighted in bold and underline respectively.
Datasets WildRF SynthWildX Average
[…]
CNNSpot (training-based) 84.20/ 82.94 79.58/ 79.27 86.62/ 86.46 91.21/ 91.85 83.72/ 83.39 87.19/ 87.26 85.42/ 85.20
UnivFD (training-based) 49.83/ 48.65 55.54/ 56.35 54.89/ 53.67 41.77/ 38.22 55.21/ 56.05 45.88/ 42.83 50.52/ 49.29
FreDect (training-based) 46.95/ 47.27 51.97/ 52.62 48.25/ 48.21 45.26/ 45.65 46.77/ 45.38 50.99/ 52.86 48.37/ 48.67
CLIP-ViT-Large-14 (zero-shot) 45.64/ 45.52 38.93/ 34.05 44.07/ 42.94 42.33/ 38.19 42.26/ 38.56 42.37/ 37.39 42.60/ 39.44
InternLM-XC.2-vl (InternLM2-7B) 42.87/ 38.76 49.93/ 51.08 42.36/ 38.97 50.06/ 52.65 56.39/ 59.33 48.39/ 49.68 48.67/ 48.41
InstructBLIP (Vicuna-7B) 53.75/ 54.82 40.54/ 34.92 45.65/ 42.31 41.24/ 36.38 42.24/ 38.63 49.56/ 48.13 45.83/ 42.53
Qwen-VL (Qwen-7B) 40.85/ 37.65 36.46/ 26.29 35.16/ 22.67 36.44/ 27.06 39.15/ 32.64 34.85/ 20.34 36.48/ 27.11
IDEFICS-Instruct (LLaMA-7B) 46.69/ 45.90 45.96/ 43.65 41.24/ 34.70 37.41/ 27.92 39.93/ 33.36 40.29/ 32.50 41.92/ 36.69
LLaVA-v1.5 (Vicuna-7B, baseline) 58.79/ 62.41 51.66/ 49.45 50.67/ 51.70 53.30/ 56.43 49.03/ 49.68 54.70/ 56.45 53.36/ 54.78
mPLUG-Owl2 (LLaMA-7B, baseline) 58.52/ 59.66 51.96/ 50.30 47.72/ 45.29 48.23/ 48.05 46.34/ 48.33 47.67/ 46.16 51.74/ 49.63
FakeScope (Ours, mPLUG-Owl2) 91.89/ 91.45 85.31/ 85.43 92.96/ 92.14 89.02/ 89.43 88.04/ 88.56 83.72/ 84.04 88.49/ 88.51
FakeScope (Ours, LLaVA-v1.5) 80.01/ 78.60 67.94/ 73.23 75.48/ 76.09 87.37/ 87.44 78.94/ 81.76 81.99/ 83.40 78.46/ 80.09
TABLE VI: Generalization capability validated on AntifakePrompt [71] and AIGCDetectBenchmark [72]. The AP (%) of various classifiers across 17 individual generators are reported. Chance is 50% and best possible performance is 100%. All the LMMs are tested under the quantitative setting. Values in gray represent the generators seen in training. No content overlapping exists in the results. The best model in each column is highlighted in bold.
Datasets AIGCDetectBenchmark AntifakePrompt Unseen Average
[…]
CNNSpot 64.86 61.41 70.33 60.10 63.08 56.41 70.19 82.85 92.51 86.50 56.61 60.25 48.88 52.15 50.54 50.22 95.15 60.71 66.00
UnivFD 61.37 50.36 39.78 55.61 48.49 51.91 56.22 43.25 35.75 44.24 45.36 58.17 50.85 57.46 80.99 60.68 68.42 57.24 53.47
FreDect 49.93 55.13 49.99 53.26 54.93 49.97 53.61 29.79 37.78 38.36 35.80 44.46 49.86 49.93 48.02 49.81 56.10 50.63 47.45
CLIP-ViT-Large-14 89.22 71.12 94.92 88.58 64.40 81.17 60.43 29.20 31.87 37.28 49.44 70.85 62.86 63.26 85.15 71.71 41.85 71.38 64.31
InternLM-XC.2-vl 52.15 46.40 61.16 50.47 52.28 44.29 68.72 35.73 41.03 48.59 45.49 62.15 52.39 58.33 60.58 58.46 71.70 58.23 53.52
InstructBLIP 50.27 40.79 32.31 88.63 40.42 49.36 55.46 25.98 29.87 25.86 72.11 64.04 59.18 74.46 95.64 81.84 97.48 67.17 57.86
Qwen-VL 72.27 42.91 41.39 82.54 44.80 48.92 84.18 75.39 69.55 72.97 83.83 88.15 66.50 78.31 93.55 89.05 91.31 73.52 72.10
IDEFICS-Instruct 40.26 51.87 53.68 39.40 47.26 66.99 35.16 22.03 23.55 25.85 22.45 41.24 46.53 41.11 34.81 44.27 48.49 45.36 40.29
mPLUG-Owl2 58.44 48.91 83.24 51.39 45.26 41.84 69.79 45.10 46.10 44.71 55.03 74.34 50.71 57.82 77.64 68.28 71.16 62.86 58.22
LLaVA-v1.5 40.97 44.09 41.30 38.08 43.03 56.01 42.36 24.32 25.07 26.71 32.78 45.13 47.93 45.08 37.69 43.91 80.87 47.40 42.08
FakeScope (mPLUG-Owl2) 89.70 92.96 88.63 89.83 86.94 93.32 98.20 97.27 99.17 98.80 99.04 99.77 98.20 98.99 99.72 98.80 99.96 95.67 95.84
FakeScope (LLaVA-v1.5) 99.92 98.01 100.00 99.66 98.01 99.45 98.96 98.21 99.04 99.10 99.04 99.65 82.59 95.97 96.40 97.20 73.64 94.68 96.17
VI-C In-the-wild Performance
To evaluate the real-world applicability of our model, we conduct in-the-wild experiments on diverse, unstructured image data collected from social media, including WildRF [73] and SynthWildX [4]. These datasets contains real-world samples with various distortions, such as compression artifacts, post-processing modifications, and noise, which pose additional challenges for forensic models. The testing protocol aligns with other experiments, and all LMMs are evaluated in the quantitative setting. The AP and AUC values for detection are listed in Table V. As reflected by the results, FakeScope equipped with mPLUG-Owl2 [37]consistently outperforms other state-of-the-art methods across the majority of generators, achieving superior overall performance in the in-the-wild testing. Notably, it demonstrates competitive results on images generated by Dalle3 and Midjourney, ranking second only to CNNSpot [20]. These findings highlight FakeScope’s robustness and effectiveness in handling unconstrained image content and generation models, reinforcing its adaptability to real-world forensic challenges beyond in-lab benchmarks.
Figure 7: Accuracy as a function of detection threshold. For each model, we present the relations between the detection threshold (x-axis) and accuracy (y-axis) on all generators of five testing databases. It turns out the FakeScope of two versions of base models (the last two sub-figures) is the only one that can achieve a great bargain between threshold sensitivity and detection accuracy, indicating remarkable robustness.
Figure 8: Performance on FakeBench [7] with respect to training data scale, demonstrating the effectiveness of FakeInstruct data. The x-axis indicates the proportion of FakeInstruct used for training and the y-axis is the corresponding performance indicator of each ability dimension.
VI-D Generalization Capability
For forensic detectors, the generalization ability on unacquainted image contents and generators is crucial for practical deployment. To assess this, we conduct a more challenging experiment to evaluate whether FakeScope is well generalizable on unseen image content and generators. Practically, we test all methods on the uncontaminated content proportions of AntifakePrompt [71] and AIGCDetectBenchmark [72], in comparison with the data employed in FakeInstruct. In particular, the AntifakePrompt dataset covers several OOD tasks beyond the primary scope of this work, such as image super-resolution, face spoofing [75], and inpainting, which pose a greater challenge to model generalization and robustness. In Table VI, the AP and AUC scores of each model across 17 image generators are presented, including 6 seen and 11 unseen ones. We also present the average AP over unseen generators as well as over all generators. From Table VI, it can be observed that the two variants of the FakeScope can achieve SOTA performance across all 17 generators and, notably, outperforms all baselines on the set of unseen generators. These results demonstrate that FakeScope can effectively handle novel image content and generation techniques, indicating strong generalization capability.
Furthermore, we analyze model stability by plotting detection accuracy as a function of the decision threshold ($\in(0,1)$) for each method, thereby assessing their sensitivity to threshold selection across different generators. In Fig. 7, a separate curve is shown for each forensic detector by averaging detecting accuracy across all generators in each dataset. Unlike AUC and mAP, the accuracy (Acc) is highly sensitive to threshold variation, and models with unstable performance may suffer under suboptimal thresholds, leading to less robustness. It can be observed from Fig. 7 that both versions of FakeScope exhibit highly consistent peak accuracy across different datasets and maintain stable performance over a broad range of threshold values. In contrast, other models show considerable variability in optimal accuracy across datasets and are overly sensitive to specific threshold settings. This indicates that FakeScope possesses prominent robustness towards forensic threshold across different datasets and generators, thereby enhancing its applicability in real-world scenarios.
VI-E Ablation Studies
To investigate the influence of different data components and scales on training FakeScope, we further conduct ablation studies on the influence of three aspects during tuning with mPLUG-Owl2 [37]: (1) scale of training data; (2) collaboration of FakeChain and visual instructions; and (3) participation of multimodal contextual information.
1) Effects of Training Data Scale: To effectively handle data-scarce scenarios is crucial for forensic models, yet it remains increasingly challenging due to the high cost and difficulty of acquiring high-quality training data. To evaluate the effectiveness of FakeInstruct dataset, we randomly sample four subsets—80%, 60%, 40%, and 20%—while preserving the original data composition, and analyze the impact on training outcomes. As illustrated in Fig. 8, the resulting FakeScope variants exhibit remarkable resilience to performance degradation as the available training data decreases, demonstrating the strong utility of FakeInstruct. Notably, the transparency-related performance remains a significant improvement over the baselines, even at the extreme 20% data proportion, and continues to strengthen as the training data size grows. This trend highlights the effectiveness of the multimodal forensic knowledge embedded in FakeInstruct, further demonstrating that the reasoning data can substantially enhance the forensic capabilities of large multimodal models.
2) Effects of Collaborative Training: During training, we combine different subsets to jointly optimize a unified forensic foundation model. To validate the effectiveness of this strategy, we compare it against task-specific training focused solely on transparent forensic capabilities. As indicated by the results listed in Table VII, the collaborative training scheme—integrating long-form reasoning with granular visual instructions—yields significant improvements in model performance. Although the various subsets in FakeInstruct originate from a shared data source, this strategy remains cost-efficient and enables FakeScope to generalize more effectively across diverse forensic tasks, demonstrating greater stability across a broader knowledge domain.
TABLE VII: Comparison on transparent forensic capabilities between only FakeChain and full FakeInstruct as training dataset. The best performance of each column is highlighted in bold.
Training Strategy Detect Interpret Reason Analyze
none (baseline) 58.22% 0.423 0.296 0.368
only FakeChain 83.27% 0.593 0.576 0.350
full FakeInstruct 99.68% 0.638 0.614 0.540
TABLE VIII: Comparison on forensic detection ability between full FakeInstruct, only absolute judgment, and only contextual data. The AP (%) on each dataset is reported. The best performance of each column is highlighted in bold.
Dataset/ Training Strategy none (baseline) only absolute only contextual full (employed)
FakeBench [7] 68.37 94.06${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+25.96}}}$ 84.13${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+15.76}}}$ 99.94${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+31.57}}}$
AntiFakePrompt [71] 60.06 90.12${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+30.06}}}$ 74.84${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+14.78}}}$ 98.90${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+38.84}}}$
AIGCDetecBenchmark [72] 54.84 81.65${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+26.81}}}$ 60.73${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+5.89}}}$ 90.23${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+35.39}}}$
WildRF [73] 52.73 82.34${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+29.61}}}$ 62.54${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+9.81}}}$ 90.05${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+37.32}}}$
SynthWildX [4] 47.41 82.78${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+35.37}}}$ 52.59${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+5.18}}}$ 86.93${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+39.52}}}$
3) Effects of Multimodal Awareness: A defining feature that sets FakeScope apart from conventional binary forensic detectors is its integration of rich multimodal contextual information on visual trace evidence during training. While these data are aimed at enhancing transparency-related forensic capabilities, they might affect the performance on simple detection. To validate whether such information also improves pure forensic detection, we compare the performance of models trained with and without it. As compared in Table VIII, incorporating contextual information not only improves the model’s transparency but also enhances detection accuracy and generalization, even when using the same amount of image data. In contrast, omitting absolute authenticity labels from training results in inconsistent improvements. These findings suggest that multimodal forensic cues can strengthen the model’s task awareness and fine-grained visual comprehension, enabling it to more effectively attend to forensic-relevant patterns. However, the model’s detection performance still heavily relies on absolute authenticity judgments, as relying solely on descriptive cues may introduce ambiguity into the decision-making process for authenticity.
VII Conclusions and Future Work
In this paper, we propose a multimodal paradigm for transparent AI-generated image forensics, versatile in detection, in-depth analysis, and extended discussions. Based on the advanced cross-modal capability of large multimodal models (LMMs), we collect and leverage rich image forensic knowledge to enhance their forensic awareness. By emulating human causal inference based on visual trace evidence, we introduce a human-machine collaborative strategy and construct the first large-scale dataset containing structured forensic reasoning data. This reasoning foundation is further expanded into a million-scale visual instruction dataset, which is internalized into LMMs to improve their transparent forensic abilities on AI-generated images. Extensive experiments demonstrate consistent effectiveness of this paradigm across LMMs with varied meta frameworks. The proposed expert forensic model can reliably handle multiple forensic tasks under a unified paradigm, and exhibits strong generalization and in-the-wild capabilities. Besides, we discovered that LMMs can effectively align with hard authenticity labels and demonstrate impressive generalization on estimating numerical probabilities, even without explicit supervision.
This work not only presents a viable pathway toward transparent AI-generated image forensics, but also offers insights into the development of trustworthy AI risk management within the broader multimodal landscape. Moreover, the proposed model could serve as the source of generating cross-modal forensic information, which may further support the development of more reliable forensic systems. Notwithstanding the use of advanced models such as GPT-4V to distill forensic knowledge, it may still represent an early-stage effort in the evolving trajectory of large multimodal models. While some uncertainties remain, this study nonetheless demonstrates the effectiveness of multimodal information and the feasibility of forensic transparency. We intend to further explore the abilities of newly emerged LMMs and contribute model update to community. Looking ahead, future efforts may focus on enhancing the model’s few-shot adaptability and dialectical reasoning capability for better robustness and explanatory power in more complex forensic contexts.

### v2

@@
 Abstract
-The rapid and unrestrained advancement of generative artificial intelligence (AI) presents a double-edged sword: while enabling unprecedented creativity, it also facilitates the generation of highly convincing deceptive content, undermining societal trust.
-As image generation techniques become increasingly sophisticated, detecting synthetic images is no longer just a binary task—it necessitates interpretable, context-aware methodologies that enhance trustworthiness and transparency.
-However, existing detection models primarily focus on classification, offering limited explanatory insights into image authenticity.
-In this work, we propose FakeScope, an expert multimodal model (LMM) tailored for AI-generated image forensics, which not only identifies AI-synthetic images with high accuracy but also provides rich, interpretable, and query-driven forensic insights.
-To this end, we first construct FakeChain dataset that contains linguistic authenticity reasoning based on visual trace evidence, developed through a novel human-machine collaborative framework.
-Building upon it, we further present FakeInstruct, the largest multimodal instruction tuning dataset containing 2 million visual instructions tailored to enhance forensic awareness in LMMs.
-Leveraging the knowledge of FakeInstruct, FakeScope achieves state-of-the-art performance in both closed-ended and open-ended forensic scenarios.
-It can distinguish synthetic images with high accuracy while offering coherent and insightful explanations, free-form discussions on fine-grained forgery attributes, and actionable enhancement strategies.
-Notably, despite being trained exclusively on qualitative hard labels, FakeScope demonstrates remarkable zero-shot quantitative capability on detection, enabled by our proposed token-based probability estimation strategy.
-Furthermore, FakeScope exhibits strong generalization and in-the-wild ability, ensuring its applicability in real-world scenarios.
+The rapid and unrestrained advancement of generative artificial intelligence (AI) presents a double-edged sword.
+While enabling unprecedented creativity, it also facilitates the generation of highly convincing content, undermining societal trust.
+As image generation techniques become increasingly sophisticated, detecting synthetic images is no longer just a binary task—it necessitates explainable methodologies to enhance trustworthiness and transparency.
+However, existing detection models primarily focus on classification, offering limited explanatory insights.
+To address these limitations, we propose FakeScope, an expert large multimodal model (LMM) tailored for AI-generated image forensics, which not only identifies synthetic images with high accuracy but also delivers rich query-contingent forensic insights.
+At the foundation of our approach is FakeChain, a large-scale dataset containing structured forensic reasoning based on visual trace evidence, constructed via a human-machine collaborative framework.
+Then we develop FakeInstruct, the largest multimodal instruction tuning dataset to date, comprising two million visual instructions that instill nuanced forensic awareness into LMMs.
+Empowered by FakeInstruct, FakeScope achieves state-of-the-art performance in both closed-ended and open-ended forensic scenarios.
+It can accurately distinguish synthetic images, provide coherent explanations, discuss fine-grained forgery artifacts, and suggest actionable enhancement strategies.
+Notably, despite being trained exclusively on qualitative hard labels, FakeScope demonstrates remarkable zero-shot quantitative detection capability via our proposed token-based probability estimation strategy.
+Furthermore, it shows robust generalization across unseen image generators and performs reliably under in-the-wild scenarios.
 The data, model, and demo will be publicly released on https://github.com/Yixuan423/FakeScope.
-Index Terms:
-AI-generated image detection, large multimodal models, forensic investigation, trustworthiness.
 I Introduction
 With the rapid development of image generation techniques, AI-synthesized images are becoming increasingly indistinguishable from genuine ones [76, 85].
-While this prosperity has driven innovation, it also introduces significant cybersecurity threats, including misinformation, propaganda, and fraud.
+While this prosperity has driven innovation, it also introduces significant cybersecurity risks, including misinformation, propaganda, and fraud.
 In response, substantial efforts have been made to detect AI-generated content in recent years [53, 12, 13, 60].
@@
 Addressing these challenges requires enhancing the transparency of AI-generated image detection models.
-We advocate for a paradigm shift in which image synthesis detection goes beyond binary classification, incorporating explainability to provide contextual insights into authenticity assisting trustworthiness.
+We advocate for the potential in which image synthesis detection goes beyond binary classification, incorporating explainability to provide contextual insights into authenticity, assisting trustworthiness.
 Figure 1: Comparative illustration of image forensic model evolution.
-(a) Binary detectors; (b) general LMMs; and (c) FakeScope, which translates visual abnormalities into human-understandable cues—much like trace evidence in forensics—using natural language.
+(a) Binary detectors; (b) general LMMs; and (c) FakeScope, which translates visual abnormalities into human-understandable cues, much like trace evidence in forensics, using natural language.
 Figure 2: Contributions of this work.
 (a) FakeChain dataset, containing reasoning on image authenticity, constructed via the proposed ACoTI strategy (Sec.
-III); (b) FakeInstruct dataset, containing 2 million visual instructions of image forensic knowledge (Sec.
+III); (b) FakeInstruct dataset, containing two million visual instructions concerning image forensic knowledge (Sec.
 IV); (c) FakeScope model, the expert model for transparent AI-generated image forensics, capable of multi-dimensional forensic capabilities (Sec.
@@
 Until recently, AI-generated image detection was predominantly framed as a binary classification task.
-Conventional approaches rely on unimodal architectures such as convolutional neural networks (CNNs) and vision transformers (ViTs) to classify images as either real or synthetic, significantly aiding in mitigating the spread of synthetic content.
+Conventional approaches rely on unimodal architectures such as convolutional neural networks (CNNs) and vision transformers (ViTs) to classify images as either real or synthetic, significantly aiding in mitigating the spread of synthetic content [68, 22, 24].
 However, these early methods largely overlooked the need for transparency, offering little more than abstract heatmaps derived from deep neural network features as visual explanations.
@@
 As illustrated in Fig.
-1 (a), different detectors can yield inconsistent decisions, significantly reducing practical utility.
-In contrast, in Fig.1(c), a transparent model can provide additional insights via visual trace evidence to enhance interpretability.
-Therefore, we anticipate more trustworthy paradigms that exceed binary detectors for achieving transparent forensics of AI-generated images.
-Such a model would enable human oversight, mitigate detection biases, and foster better confidence in AI-generated content forensics [2].
-With the advent of multimodal models, visual-language approaches have started to emerge.
+1(a), different detectors can yield inconsistent decisions, significantly reducing practical utility.
+In contrast, in Fig.1(c), a transparent [^1] Following [28, 44], transparency herein refers to a composite spectrum including explainability, traceability, and justifiability.
+forensic model can go beyond binary decisions by offering visual-evidence-based justifications, supporting query-driven forensic analysis, and estimating authenticity probabilities in a unified and interpretable manner.
+Such capabilities are key to enabling human oversight, mitigate detection biases, and foster better confidence in AI-generated content forensics [28, 85].
+With the advent of multimodal models, vision-language approaches have started to emerge [66, 31, 80, 9, 29, 71, 84].
 Zhang et al.
@@
 [77] explored the capabilities of large multimodal models (LMMs) in the forensic analysis of AI-generated images.
-In [11, 12, 13], LMMs are incorporated to localize tampering forgeries and exhibits substantial effectiveness.
+In [75, 46, 32], LMMs are incorporated to localize tampering forgeries and exhibit substantial effectiveness.
 Besides, LMMs are employed to assist in improving detection accuracy.
@@
 Wen et al.
-[16] train a binary classifier with ViT intermediate features and use LMMs to explain the detection.
-These studies establish a concrete foundation for the transition from unimodal to multimodal approaches in the area of AI-generated image forensics.
-Nevertheless, due to the inherent limitations of LMMs as classifiers [17], previous approaches have struggled to balance detection accuracy and transparency, often relying on separate strategies and training procedures to fulfill multi-dimensional forensic tasks.
-Therefore, there is an urgent need for a unified forensic expert model for AI-generated images capable of multiple forensic tasks in diversified scenarios.
-Inspired by the above insights, we present FakeScope, a unified expert model for transparent and versatile AI-generated image forensics.
-It leverages the remarkable cross-modality ability of generally pretrained LMMs while being systematically enhanced for universal forensic awareness.
-Specifically, it is designed with three key objectives: (1) comprehension on trace evidence and authenticity; (2) faithful response to diverse user queries; and (3) qualitative and quantitative detection without model extension.
-To achieve these, we collect and leverage forensic knowledge to improve LMMs’ forensic awareness on AI-generated images.
+[71] train a binary classifier with ViT intermediate features and use LMMs to explain the detection result.
+These studies have established a concrete foundation for transitioning from unimodal to multimodal approaches in AI-generated image forensics.
+Nevertheless, due to the inherent limitations of LMMs as classifiers [81], previous approaches have struggled to achieve high detection performance and transparency simultaneously.
+Most existing works rely on separate strategies and training procedures, such as introducing additional classifier heads or independently trained explanation modules, to fulfill different forensic demands.
+Such fragmented pipeline and weak coupling can lead to disconnected probabilistic reasoning space across tasks and hindering the model from reusing shared forensic knowledge, further constraining the model’s potential.
+This motivates the need for a unified forensic expert model for AI-generated images capable of performing multiple forensic tasks in diversified scenarios.
+Inspired by the insights, we present FakeScope, a unified expert model for transparent and versatile AI-generated image forensics.
+It leverages the remarkable cross-modality ability of generally pretrained LMMs while being systematically enhanced for forensic awareness.
+Specifically, it is designed with three key objectives: (1) comprehension of visual trace evidence and authenticity; (2) faithful response to diverse user queries; and (3) qualitative and quantitative detection without model extension.
+To achieve these, we collect and leverage domain-specific knowledge to improve LMMs’ forensic awareness of AI-generated images.
+Our emphasis is therefore on output-level image forensic transparency rather than on modeling the causal interpretability of internal LMM mechanisms.
 As shown in Fig.
-2(a), we first construct FakeChain, a large-scale multimodal dataset containing long and comprehensive reasoning on image authenticity as forensic knowledge source, developed via a cost-efficient human-in-the-loop strategy.
-We further enrich FakeChain to create FakeInstruct, a gigantic-scale dataset comprising 2 million visual instructions to enhance diverse forensic abilities of LMMs (Fig.
+2(a), we first construct FakeChain, a large-scale multimodal dataset containing long and comprehensive reasoning on image authenticity as a forensic knowledge source, developed via a cost-efficient human-in-the-loop strategy.
+We further enrich FakeChain to create FakeInstruct, a dataset comprising two million visual instructions to enhance the diverse forensic abilities of LMMs (Fig.
 2(b)).
@@
 2(c)) exhibits state-of-the-art (SOTA) performance in detecting, explaining, analyzing, and discussing AI-generated images.
-In particular, it can generalize well to unseen image content, generation models, and even user-created contents.
+In particular, it generalizes well to unseen image content, generation models, and user-created content.
 Our core contributions are as follows.
-  • We present FakeChain, a large-scale multimodal dataset with coherent reasoning flow from visual trace evidence to image authenticity.
-It is collected through our proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) scheme, a cost-effective human-machine collaboration strategy that requires minimal human annotation while ensuring reliability.
-  • We develop FakeInstruct, the first million-scale dataset designed to enhance LMMs’ multi-dimensional forensic capabilities in AI-generated image detection.
-  • We propose FakeScope, a multimodal expert model for transparent AI-generated image detection that integrates both detection and query-based forensic analysis.
-We further propose a zero-shot token-based strategy for estimating detection probabilities.
-Extensive experiments validate FakeScope’s superiority in both detection and transparency-related tasks, outperforming other LMMs and binary models across diverse datasets.
-II Related Works
-Recently, the rapid advancement of image generation technology has driven growing interest in AI-generated image forensics, and the evolution of LMMs has opened new avenues for enhancing the trustworthiness of image forensic systems.
-We first outline the progression of AI-generated image detection techniques, followed by an overview of the current developments in LMMs.
+  • We present FakeChain, a large-scale multimodal dataset with structured reasoning from visual trace evidence to image authenticity.
+It is built with our proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) scheme.
+This cost-effective human-machine collaboration strategy requires minimal human annotation while ensuring data reliability.
+  • We develop FakeInstruct, the first million-scale visual instruction dataset for enhancing LMMs’ forensic capabilities in AI-generated image detection, exhibiting consistent enhancement to diverse base models.
+  • We propose FakeScope, a multimodal forensic expert built on FakeInstruct for transparent AI-generated image detection.
+It unifies detection and query-based forensic analysis, and further incorporates a zero-shot token-based strategy for probability estimation.
+Extensive experiments validate its superiority in detection and transparency-related tasks, outperforming other LMMs and binary models across diverse datasets.
+II Related Work
+In this section, we first outline the progression of AI-generated image detection techniques, followed by an overview of recent developments in LMMs.
 II-A AI-generated Image Detection
@@
 This problem is commonly formulated as a binary classification task.
-Preexistent approaches generally incorporate forensic patterns to classify, which can be categorized into the flaw-based [18, 19] and fingerprints-based [20, 21, 22, 23] according to their forensic ground.
-The flaw-based methods capture the intricate details in low (e.g., texture [18]) or high (e.g., physical rules [19, 24]) semantic levels, while the fingerprints-based detectors identify the invisible leftover patterns of generators (e.g., frequency artifacts [20, 21, 22, 23]).
+Preexisting approaches generally incorporate forensic patterns to classify, which can be categorized into the flaw-based [83, 11] and fingerprints-based [68, 22, 24, 72] according to their forensic ground.
+The flaw-based methods capture the intricate details in low (e.g., texture [83]) or high (e.g., physical rules [11, 63]) semantic levels.
+At the same time, the fingerprints-based detectors identify the invisible leftover patterns of generators (e.g., frequency artifacts [68, 22, 24, 72]).
 Moreover, investigations on model’s generalization ability towards unseen  [43, 33, 12, 53, 13, 60] and generator attribution [64, 69] have emerged.
-Some of them assume to be only accessible to real images and utilize generators (like an autoencoder) to simulate generative artifacts [29, 22, 26, 25].
-Although remarkable, current fake image detection models still fall short in the interpretability towards human [30], which causes the detection results hard-to-explain for the general populace and weakens the credibility of model output.
-Therefore, developing an expert dedicated to fake image detection with both high detection accuracy and transparency is currently urgently needed.
+Some of them are assumed to be only accessible to real images and utilize generators (like an autoencoder) to simulate generative artifacts [79, 24, 33, 43].
+Although remarkable, current fake image detection models still fall short in interpretability towards humans [41], which makes the detection results hard to explain for the general populace and weakens the credibility of the model output.
+Therefore, developing an expert dedicated to fake image detection with high detection accuracy and transparency is urgently needed.
 II-B Large Multimodal Models
-Typical LMMs compromise a modality encoder, a large language model (LLM), with a modality interface for cross-modal connection [31, 32].
-The specialty of LMM lies in their ability to effectively receive, process, and generate multimodal information.
-Currently, LMMs are developing by leaps and bounds and have exhibited remarkable visual-language capabilities on par with or even surpassing the human, including the closed-source models (e.g., GPT-4V [33], Gemini [34], and Claude [35]) and the open-source ones (e.g., LLaVA [36], mPLUG-Owl [37], deepseek [38]).
+Typical LMMs comprise a modality encoder, a large language model (LLM), with a modality interface for cross-modal connection [20, 30].
+The specialty of LMMs lies in their ability to effectively receive, process, and generate multimodal information.
+Currently, LMMs are developing by leaps and bounds and have exhibited remarkable visual-language capabilities on par with or even surpassing the human, including the closed-source models (e.g., GPT-4 [54], Gemini [67], and Claude [3]) and the open-source ones (e.g., LLaVA [45], mPLUG-Owl [78], Deepseek [49]).
 Furthermore, to adapt general-purpose LMMs to specialized fields, recent researchers proposed various expert models based on visual instruction tuning  [48, 50, 21, 45].
-Despite great progress, recent studies [7] indicate that LMMs’ forensic capabilities are still unsatisfactory, especially when it comes to identifying fine-grained forgery aspects.
-On the other hand, human crowd-sourcing data for visual instruction tuning becomes increasingly hard to get, partially due to the process being time-consuming, especially when multimodal information is considered.
-In this work, to infuse the forensic abilities on AI-generated images into foundational LMMs, we construct the first visual instruction tuning dataset centering on fake image detection, the FakeInstruct, aiming at a unified multimodal paradigm based on LMMs for more transparent image forensics.
+Despite great progress, recent studies [40] indicate that LMMs’ forensic capabilities are still unsatisfactory, especially when identifying fine-grained forgery aspects.
+In this work, to infuse the forensic abilities of AI-generated images into foundational LMMs, we construct the first visual instruction tuning dataset centering on fake image detection, the FakeInstruct, aiming at a unified multimodal paradigm based on LMMs for a more transparent image forensics.
 III FakeChain: Reasoning on Image Authenticity
-In this section, we elaborate on the construction of FakeChain, a large-scale dataset with long and comprehensive reasoning on underlying trace evidence for authenticity.
-It functions as the source of forensic knowledge and further contributes to building the instruction dataset and the forensic model.
-For construction, we first prepare images (Sec.
+In this section, we elaborate on the construction of FakeChain, a large-scale dataset of visual trace-based authenticity reasoning.
+FakeChain serves as the primary source of forensic knowledge and further supports the construction of the instruction dataset and the expert model.
+We first prepare images (Sec.
 III-A) and then collect authenticity reasoning.
-While human annotation remains the most intuitive and widely applied approach for training LMMs [42, 36], acquiring sufficiently fine-grained dataset for authenticity reasoning is highly labor-intensive.
-To mitigate the prohibitive costs, a natural idea is to use a human-machine collaboration strategy.
-Therefore, we first collect human expert feedback (Sec.
-III-B) as weak supervision in the proposed Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy (Sec.
-III-C) for forensic reasoning collection.
+While human annotation remains the most intuitive approach for training LMMs [73, 45], collecting fine-grained authenticity reasoning at scale is highly labor-intensive.
+To reduce this cost, we adopt a human-machine collaborative strategy (Fig.
+3) that uses limited human annotations as weak supervision in three steps:
+1. Steer, which guides reasoning with human-annotated trace-evidence categories;
+2. Demonstrate, which provides human reasoning exemplars;
+3. Enlighten which elicits visual reasoning from strong models through few-shot chain-of-thought prompting.
+Building on this scheme, we first collect human expert feedback (Sec.
+III-B) and propose Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy (Sec.
+III-C2) to enlighten strong models under the steer of a few human annotations assisted by human reasoning demonstrations.
+With ACoTI, we construct the FakeChain dataset.
+Details are elaborated below.
 Figure 3: Pipeline of the proposed ACoTI strategy for obtaining the reasoning on trace evidence under the human-machine collaboration scheme.
@@
 III-A Image Preparation
-FakeChain contains both AI-generated and real images in equal quantities.
-In contrast to the one-to-all generalization setting in some works [25, 26, 4, 3, 5, 6], FakeChain is deliberately constructed with a high degree of diversity to capture a broad range of generative patterns.
-In particular, AI-generated images function as the knowledge source of forensic trace evidence.
-To ensure content diversity, we sampled 23,797 AI-generated images from multiple sources.
-These generators span 17 types, including GAN-based, diffusion-based, auto-regressive, and proprietary models.
-In specific, we choose representative ProGAN [43], StyleGAN [44], BigGAN [45], CogView2 [46], ADM [47], IF [48], FuseDream [49], VQDM [50], Glide [51], SD [52], SDXL [53], FLUX.1 [54], Firefly [55], Dalle2 [56], Dalle3 [57], Wukong [58] and Midjourney [59].
-Also, we collect 23,797 real images from DIV2K [60], RAISE [61], and ImageNet [62].
-The details of image components are provided in the supplementary material (SM).
+FakeChain contains a equal number of AI-generated images and real images, totaling 47,594.
+Unlike the one-to-all generalization setting in [43, 33, 12, 53, 13, 60], FakeChain is deliberately designed for high diversity to cover a wide range of generative patterns.
+In particular, AI-generated images function as the knowledge source of forensic trace evidence and are sampled from multiple sources.
+The generators span 17 types, including GAN-based, diffusion-based, auto-regressive, and proprietary models.
+Specifically, we choose representative ProGAN [35], StyleGAN [36], BigGAN [6], CogView2 [19], ADM [18], IF [16], FuseDream [47], VQDM [27], Glide [52], SD [61], SDXL [56], FLUX.1 [23], Firefly [1], Dalle2 [58], Dalle3 [65], Wukong [74] and Midjourney [51].
+Also, we collect 23,797 real images from DIV2K [2], RAISE [15], and ImageNet [17], ensuring content diversity.
+The details are provided in the supplementary material (SM).
 III-B Human Expert Feedback
 As shown in Fig.
-3, we require a few human annotations in step1-steer and step2-demonstrate of ACoTI, including the trace evidence category and reasoning demonstration respectively.
-Therefore, we conduct a well-controlled in-lab experiment involving 30 trained human experts with experience in image generation as annotators, ensuring qualified understanding of generative patterns.
-For step1-steer, experts are required to correctly distinguish whether the image is AI-generated, and then select supporting categories, ensuring only accurate judgments contribute to the dataset.
-In practice, we provide a category set for reference $\mathcal{F}$ = {texture, edge, clarity, light&shadow, anatomy, layout, symmetry, reflection, perspective, physics, shape, theme, content deficiency, distortion, unrealistic, overall hue[7], which can characterize authenticity across different levels of trace evidence.
-Each image is annotated by two and only the intersection of selected categories is retained to improve reliability.
-Additionally, experts can provide supplementary authenticity-related information at will, such as specific area of occurrence and commonsense knowledge.
-For step 2-demonstrate, we fix 50 fake and 50 genuine images, and employ experts to annotate manually with detailed reasoning.
-These expert-annotated images constitute the exemplar sets $\{{\chi}^{+},{\chi}^{-}\}$ as illustrated in upper Fig.
+3, we require a few human annotations in step1-steer and step2-demonstrate, including the trace evidence category and reasoning demonstration.
+Therefore, we conduct a controlled in-lab experiment involving 30 trained human experts experienced in image generation, ensuring qualified understanding of generative patterns.
+In step1-steer, experts first distinguish whether an image is AI-generated and then select supporting categories.
+This ensures only correct annotations contribute to the dataset.
+In practice, we provide a category set for reference $\mathcal{F}$ = {texture, edge, clarity, light&shadow, anatomy, layout, symmetry, reflection, perspective, physics, shape, theme, content deficiency, distortion, unrealistic, overall hue[40], covering multiple levels of forensic evidence.
+To avoid overly broad selections, annotators are instructed to choose only verifiable cues rather than all plausible categories.
+Practically, each image receives at most six categories.
+Every image is annotated by two experts, and only the intersection of their selections is kept to reduce annotator-specific bias.
+Additionally, experts can provide authenticity-related observations at will, such as specific affected regions or relevant commonsense cues.
+In step 2-demonstrate, we fix 50 fake and 50 genuine images and employ experts to annotate manually with detailed reasoning.
+These annotations constitute the exemplar set $\{{\chi}^{+},{\chi}^{-}\}$ shown in upper Fig.
 3, where + and - denote AI-generated and genuine images respectively.
-To ensure balance, we manually guarantee each category of $\mathcal{F}$ emerges in at least three exemplars.
-Herein, we establish the necessary human supervision for the subsequent reasoning collection.
+To ensure coverage, we manually guarantee that each category in $\mathcal{F}$ emerges in at least three exemplars.
+This process establishes the human supervision needed for subsequent reasoning collection.
+Figure 4: The composition of the FakeInstruct, which is derived from the FakeChain dataset, containing 47K visual instructions on absolute authenticity judgments, 95K instructions on holistic reasoning, 715K on fine-grained visual trace evidence, and 1190K on extended knowledge and discussions.
+The million-scale diversified visual instructions of FakeInstruct enable LMMs with a broad and fine-grained understanding of image authenticity, ensuring LMMs are equipped to handle diverse forensic tasks with both interpretability and accuracy.
 III-C Reasoning on Visual Trace Evidence
 III-C1 Task Definition
-FakeChain encapsulates forensic knowledge to enable models a fundamental understanding between visual trace evidence and authenticity.
-We structure it as a materialized reasoning process, comprising an exhaustive description of the appearance and location of trace evidence, enclosed with an authenticity conclusion.
-Compared with free-form descriptions, this structured reasoning format has two advantages: (1) it allows for a precise and detailed reflection of the aspects humans focus on when assessing authenticity; and (2) it inherently follows the logical pattern of human thought.
-For instance, when examining a suspicious image, humans first scrutinize telltale visual clues and then summarize an inclined conclusion like “Thus, I tend to think this is a fake image generated by AI”.
-This structured reasoning approach helps LMMs more effectively mimic human decision-making in AI-generated image detection, thereby enhancing human-centered transparency.
-Figure 4: The composition of the FakeInstruct which is derived from the FakeChain dataset, containing 47K visual instructions on absolute authenticity judgments, 95K instructions on holistic reasoning, 715K on fine-grained visual trace evidence, and 1190K on extended knowledge and discussions.
-The million-scale diversified visual instructions of FakeInstruct enable LMMs with both broad and fine-grained understanding of image authenticity, ensuring LMMs are equipped to handle diverse forensic tasks with both interpretability and accuracy.
+FakeChain encodes forensic knowledge as a materialized reasoning process that links exhaustive description of the appearance and location of trace evidence to an authenticity conclusion.
+Compared with free-form descriptions, this structured format has two advantages: (1) it allows for a precise and detailed reflection of the aspects humans focus on when assessing authenticity; and (2) it follows the pattern of human decision-making.
+By organizing supervision in this way, FakeChain can help to improve human-centered transparency.
+We emphasize, however, that this materialized reasoning should not be equated with the actual causal basis underlying the model’s authenticity judgment, but rather as a post-hoc explanation.
 III-C2 ACoTI Strategy
-To develop a cost-efficient approach for obtaining large-scale reasoning data, we present a semi-automatic scheme to efficiently extract forensic knowledge from AI experts with few human annotations as supervision, i.e., the Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy.
+To obtain large-scale reasoning data cost-effectively, we propose a semi-automatic scheme for extracting forensic knowledge from AI experts with marginal human supervision: Anthropomorphic Chain-of-Thought Inference (ACoTI) strategy.
 As illustrated in Fig.
-3, ACoTI operates in a human-machine collaborative framework, where humans partially participate in the first and second phases by providing partial information of trace evidence.
-Inspired by human reasoning processes, ACoTI follows the Steer-Demonstrate-Enlighten pipeline, with the entire process detailed as follows.
+3, ACoTI follows a human-machine collaborative framework in which humans participate only in the first two stages by providing partial trace-evidence information.
+This method follows a Steer-Demonstrate-Enlighten pipeline, detailed as follows.
 Step 1.
-Steer: To utilize off-the-shelf LMMs maximally on reasoning, we employ the category of visual trace evidence to steer the reasoning process.
-We first notify the AI expert with the human-annotated category set $\mathcal{F}_{i}$ for the $i^{th}$ target image to ensure reasoning process focused on forensic trace evidence rather than drifting into unrelated domains.
-The $\mathcal{F}_{i}$ directly participates in the following two steps.
+Steer: For the $i$-th candidate image, we use the human-annotated trace-evidence category set $\mathcal{F}_{i}$ to steer the reasoning process.
+In practice, $\mathcal{F}_{i}$ is carried into the next two stages so that reasoning remains grounded in forensic evidence rather than drifting to irrelevant content.
 Step 2.
-Demonstrate: To further mitigate hallucinations in imperfect models, we leverage human demonstrations as in-context examples, inspired by LMMs’ ability to learn from explanations in context [63].
-Considering the trade-off between context length and efficiency, we adaptively select the most relevant demonstrations by filtering a minimum covering set [64] ${\chi}_{i}$ from $\{{\chi}^{+},{\chi}^{-}\}$ for $\mathcal{F}_{i}$.
-We ensure at least one positive sample and one negative sample are included in ${\chi}_{i}$ via optimizing the following process:
+Demonstrate: To further reduce hallucination, we provide human-written demonstrations as in-context examples,  [38].
+Considering the trade-off between context length and efficiency, we adaptively select a minimum covering set [5] ${\chi}_{i}$ from the exemplar set $\{{\chi}^{+},{\chi}^{-}\}$ according to $\mathcal{F}{i}$ of the target image.
+Formally, we solve the following optimization:
 $\chi_{i}=\arg\min_{\chi^{\prime}\subseteq\{\chi^{+}\cup\chi^{-}\}}|\chi^{\prime}|,$ (1)
-[…]
-where $L(\cdot)$ represents the mapping function that extracts the categories covered by a given subset of exemplars.
-This ensures the selected examples can comprehensively represent the key forensic attributes required for reasoning.
+$\text{{s.t.}}\quad L(\chi^{\prime})=\mathcal{F}_{i},\ \chi^{\prime}\cap\chi^{+}\neq\emptyset,\ \chi^{\prime}\cap\chi^{-}\neq\emptyset,$
+where $L(\cdot)$ maps a subset of exemplars to the union of trace categories they cover.
+In practice, we select 3 to 5 demonstrations from ${\chi}^{+},{\chi}^{-}$} for each target image to satisfy both coverage and label-diversity constraints.
+The selection is implemented with a greedy category-covering algorithm with random tie-breaking, yielding a compact yet diverse demonstration set that captures the key forensic attributes while remaining within the context window.
 Step 3.
-Enlighten: The core of ACoTI lies in enlightening the AI expert through a carefully structured prompting scheme.
+Enlighten: The core of ACoTI is a structured prompting scheme that guides the teacher model GPT-4V [54] to produce forensic reasoning.
 As illustrated in Fig.
-3, the proposed ACoTI prompting emulates human reasoning process with few-shot chain-of-thought (CoT) prompting [65].
-A unique aspect is explicitly providing the AI with the correct authenticity label in the context, guiding the model toward the appropriate answer space [65].
-Finally, this framework is structured as a quadruple: <context, target, human supervisions, CoT prompt>, where human supervision contains $\mathcal{F}_{i}$, $\chi_{i}$ and $auth_{i}$.
-Detailed prompts are provided in the SM.
-In this way, ACoTI provides an effective pipeline for distilling forensic knowledge from strong models into usable reasoning data.
-Compared with manually collecting long and complex descriptions from human annotators, ACoTI significantly reduces the burden while maintaining data quality.
-To quantify its effectiveness, we further conducted a two-alternative forced choice (2AFC) study on human preference between the outcomes generated with ACoTI and those without.
-Six annotators rated on 100 randomly sampled annotated images, and ACoTI demonstrated 99.17% preference, indicating its effectiveness as a cost-efficient approach for extracting forensic knowledge from imperfect models.
-Therefore, we apply ACoTI to all collected images and organize the resulting data into the predefined reasoning format $rsn$.
-As shown in Fig.
-4(b), we provide an image with corresponding $rsn$.
-The 47,594 tuples of $\{\bm{\mathtt{I}},auth,rsn\}$ compose the FakeChain dataset, where $\bm{\mathtt{I}}$ and $auth$ denote the image and authenticity label respectively.
-IV FakeInstruct: Million-scale Multimodal Forensic Instruction Dataset
-Although FakeChain offers corpus-rich reasoning with abundant forensic knowledge, we transform it into a diverse set of visual instructions to better leverage the information embedded in long-form text.
+3, we adopt few-shot chain-of-thought prompting[70] augmented with human supervision.
+For each target image, we construct a prompt as a quadruple <context, target, human supervision, CoT prompt> , where human supervision includes the attribute set $\mathcal{F}_{i}$, selected demonstrations $\chi_{i}$, and ground-truth label $auth_{i}$.
+A unique design choice is to explicitly provide the authenticity label in the context to anchor the reasoning toward the appropriate answer space [70].
+For example, when the attribute set $\mathcal{F}_{i}$ = {texture, shadow} and ground-truth label $auth_{i}$ = fake, the model is prompted with task context, target, and human supervision containing attribute guidance and filtered demonstrations.
+The exact prompting structure is provided in the SM.
+In this way, ACoTI provides an effective pipeline for distilling forensic knowledge from strong models into structured reasoning data.
+Compared with manually collecting descriptions from human annotators, ACoTI substantially reduces the burden while maintaining data quality.
+To quantify its effectiveness, we conducted a two-alternative forced choice (2AFC) study on human preference between the outcomes generated with ACoTI against vanilla prompting.
+Six annotators compared 100 randomly sampled images, and ACoTI outputs are preferred in 99.17% of cases, validating it as a cost-efficient strategy for extracting forensic knowledge from imperfect models.
+Based on this, we apply ACoTI to all collected images and organize the resulting data into the predefined reasoning format rsnrsn.
 As illustrated in Fig.
-2(b), we further derive instruction-following data from FakeChain in the visual question answering (VQA) format with both Mistral [66] and GPT-4o [33] in an intermingling manner to enhance linguistic and instructional diversity.
-We take extra precautions to remove textual contamination allowing models to infer answers without visual information, thereby ensuring the integrity of instruction-based reasoning [67].
-As depicted in Fig.
-4, the dataset contains 2000K diversified instructions in total, which endows LMMs with flexibility and knowledge in handling extensive forensic scenarios.
-The details are elaborated as follows.
+3, each image is paired with its rsnrsn, forming the final dataset.
+In total, the resulting FakeChain dataset consists of 47,594 tuples of $\{\bm{\mathtt{I}},auth,rsn\}$, where $\bm{\mathtt{I}}$ and authauth denote the image and its authenticity label, respectively.
+IV FakeInstruct: Multimodal Forensic Instruction Dataset
+We transform FakeChain into diverse visual instructions to exploit the information embedded in long-form text.
+As illustrated in Fig.
+2(b), we derive instruction-following data with Mistral [34] and GPT-4o [54] in an intermingling manner to enhance linguistic diversity.
+We take extra precautions to remove textual contamination, prohibiting models from answering without visual information to ensure visual reasoning [25].
+As illustrated in Fig.
+4, the dataset contains 2M visual instructions, equipping LMMs to handle a wide range of forensic scenarios.
+Details are given below.
 Part 1.
 Absolute Authenticity Judgments: As shown in Fig.
-4(a), we design diverse question formulations on absolute image authenticity to explicitly implant the concepts of fake and real in LMMs.
-Throughout this part, we utilize the term fake to denote the AI-generated images, while real for the opposite.
-While incorporating synonyms such as genuine, fault, and artificial can enhance interaction flexibility with human users, we prioritize a higher level of certainty in objective authenticity assessment in this subset.
+4(a), we design diverse questions on absolute image authenticity to explicitly instill the concepts of fake and real in LMMs.
+We use fake to denote the AI-generated images, while real for the opposite.
+We prioritize unambiguous authenticity judgments in this subset and thus not include synonyms.
+In total, this part comprises about 47K instructions.
 Part 2.
 Holistic Reasoning: As shown in Fig.
-4(b), the holistic reasoning imitates human inference on image authenticity, which can foster LMMs’ awareness of the relationships between multi-level visual trace evidence and the overall authenticity.
-Additionally, the carefully structured cause-and-effect reasoning framework enhances LMMs’ ability to distinguish AI-generated images by activating their latent reasoning capabilities [68].
+4(b), holistic reasoning induces inference on image authenticity to help LMMs connect multi-level visual trace evidence with overall image authenticity.
+The structured cause-and-effect format also helps activate latent reasoning abilities for distinguishing AI-generated images [37].This subset includes 95K instructions.
 Part 3.
-Fine-grained Trace Evidence: To equip LMMs with the ability to seamlessly respond to diverse forensic queries from users, we decompose and reformat the reasoning data in FakeChain into a variety of visual instructions, structured in narrative or multiple-choice question (MCQ) [69] formats.
-This transformation is facilitated by GPT-4o [33] and Mistral [66], whose complementary strengths ensure a diverse distribution of instructional corpus.
+Fine-grained Trace Evidence: To help LMMs handle diverse forensic queries, we decompose FakeChain into visual instructions in both narrative and multiple-choice (MCQ) formats.
+This transformation is facilitated by GPT-4o [54] and Mistral [34], whose complementary strengths improve instructional diversity.
 As illustrated in Fig.
-4(c), the dataset includes authenticity-related VQAs covering what, how, yes-or-no, and imperative question types.
-These questions pertain to perceptual adjectives (e.g., unnatural, misaligned), questionable visual attributes (e.g., smudge, shape), regions of occurrence (e.g., right eye, apple skin), or contextual knowledge (e.g., shadow in wrong directions).
-The MCQs comprise one correct answer alongside multiple plausible yet misleading distractors, thereby enhancing models’ ability to discern subtle visual cues while mitigating misconceptions.
+4(c), this subset covers authenticity-related questions in what, how, yes-or-no, and imperative forms, targeting perceptual adjectives, questionable visual attributes (e.g., shape), regions of occurrence (e.g., right eye), or contextual knowledge (e.g., shadows in the wrong direction).
+Each MCQ comprises one correct answer and several plausible misleading distractors, encouraging sensitivity to subtle cues while reducing misconceptions.
+This part contains about 715K instructions.
 Part 4.
-Extensional Knowledge: Beyond fundamental trace evidence reasoning and forensic analysis, we further incorporate extensional knowledge through open-ended discussions (Fig.
+Extensional Knowledge: Beyond forensic analysis, we further introduce extensional knowledge through open-ended discussions (Fig.
 4(d)).
-Specifically, these discussions cover four major aspects: origins of perceptual misjudgments arising from trace evidence, suggesting improvements to enhance the realism of generation, inferring potential generator characteristics based on forensic evidence, and exploring additional user-driven inquiries related to authenticity.
-We herein present a brief analysis of the FakeInstruct dataset.
-In total, about 2000K visual instructions are included in the FakeInstruct dataset.
+These discussions cover four aspects: origins of perceptual misjudgments arising from trace evidence, suggesting improvements to enhance the realism of generation, inferring potential generator characteristics, and other user-driven inquiries related to authenticity.
+This component consists of more than 1190K visual instructions.
+Finally, about 2M visual instructions are included in the FakeInstruct dataset.
 As shown in Fig.
-4(e), quantitatively, the image content type and instruction type are well-balanced.
+4(e), the image content type and instruction type are quantitatively well-balanced.
 Qualitatively, Fig.
@@
 Moreover, we propose a non-trivial probability estimation strategy via token soft-scoring, which could extend the model’s qualitative authenticity judgments to quantifiable probability estimations, without requiring additional explicit training on numerical labels.
+This probability estimation operates purely at inference as designed, and does not rely on any auxiliary trainable modules.
+We aim to enable FakeScope to support user queries beyond binary detection and to provide post-hoc, human-facing rationales that improve the accessibility of its decisions.
 The details are elaborated as follows.
 V-A Architecture and Supervised Instruction Tuning
-The architecture of FakeScope follows the LLaVA-style [36] vision-language model (VLM) framework, compromising an image encoder that transforms images into tokenized representations, an LLM for textual interpretation and generation, and a vision-language connector that bridge modalities.
-As listed in Table I, to validate the universality of FakeInstruct, we incorporate two different model architectures, including the 7B scale LLaVA-v1.5 [36] and mPLUG-Owl2 [37], corresponding to two variants of FakeScope.
-LMM training typically involves two phases: the cross-modal alignment requiring gigantic-scale data and visual instruction tuning on data of proprietary domain [36].
-Given the extensive scale of FakeInstruct, we introduce an auxiliary training stage to perform supervised visual instruction tuning specifically on FakeInstruct.
-Verified by previous studies [70], adequate tuning on the visual instructions in FakeInstruct is expected to facilitate the model’s visual forensic capabilities while also preserving the general knowledge obtained from earlier-phase training.
-To optimize efficiency, we freeze the vision encoder and fine-tune only the modality projector and language model.
+FakeScope follows the LLaVA-style vision-language architecture [45].
+To verify the generality of FakeInstruct, we instantiate FakeScope with two 7B backbones as variants, including LLaVA-v1.5 [45] and mPLUG-Owl2 [78].
+Given the scale of FakeInstruct, we further conduct supervised visual instruction tuning on this dataset to adapt the pretrained LMMs to image forensic while also preserving the general knowledge obtained from earlier-phase trainings [45].
+We freeze the vision encoder and fine-tune only the modality projector and language model for efficiency.
 V-B Probability Estimation via Token Soft-Scoring
+LMMs typically convey their forensic judgment through selecting the token with the highest logit, which inherently limits their ability to provide numerical probability scores.
+Moreover, since general-purpose LMMs are not explicitly trained as classifiers with numerical labels, direct probability prediction is not a natural formulation.
+To address this, FakeScope is trained using binary authenticity judgments and incorporates a token soft-scoring strategy that operates purely at inference time to convert model outputs into quantitative probability estimates.
+As depicted in Fig.
+5, we define fake and real as anchor tokens and apply softmax normalization over target logits within a standardized prompting framework:
+USER: <Image> What is the authenticity of this image?
+ASSISTANT: This image is <auth_token>.
+Probabilities are computed from the logits at the position of <auth_token>, where the candidate token sets are defined as $\mathcal{T}=\{\tilde{t}_{j}|_{j=0}^{1}\}=\{\tilde{t}_{fake},\ \tilde{t}_{real}\}$, where $\tilde{t}_{(\cdot)}$ represents case variations of the anchor tokens for robustness.
+The probability $p_{i}$ of each class is then calculated as follows,
+$p_{j}=\frac{e^{\sum_{v\in\tilde{t}_{j}}z^{(v)}}}{\sum_{j=0}^{1}e^{\sum_{v\in\tilde{t}_{j}}z^{(v)}}},$ (2)
+where $z^{(v)}$ represents the logit associated with token $v$.
+Unlike hard token selection, this approach retains nuanced confidence signals that would otherwise be discarded during deterministic text generation.
+This enables FakeScope to generate quantifiable probability estimates, improving its explainability as a forensic detector.
 Figure 5: Probability estimation in FakeScope.
@@
 The illustration is based on the mPLUG-Owl2 [78] backbone.
-TABLE I: Meta structures of LMMs as FakeScope foundation.
-[…]
-LLaVA-v1.5 (7B) [36] CLIP-ViT-L14 MLP Vicuna-7B
-mPLUG-Owl2 [37] CLIP-ViT-L14 Abstractor LLaMA2-7B
-TABLE II: Publicly available datasets used for evaluation. The ‡ notion indicates the uncontaminated content of the dataset, and # denotes the unit number.
+TABLE I: Publicly available datasets used for evaluation. The ‡ notion indicates the uncontaminated content of the dataset, and # denotes the unit number.
 Purpose Datasets Modality # Real/ # Fake Real Image Source # Generators Evaluation Criteria
-Multimodal Forensic Capability FakeBench [7] Multiple 3k / 3k ImageNet, DIV2K, COCO 10 Detecting, Interpreting, Reasoning, Fine-grained Analyzing
+Multimodal Forensic Capability FakeBench [40] Multiple 3k / 3k ImageNet, DIV2K, COCO 10 Detecting, Interpreting, Reasoning, Fine-grained Analyzing, Forensic Attributing
 Generalization Capability AntiFakePrompt [8] Single 42k / 6k COCO 11 Detecting
@@
 SynthWildX [12] 500 / 1.5k X 3
-LMMs typically convey their forensic judgment through hard token selection, i.e., choosing the token with the highest logit when generating responses.
-Here, “token” refers to the vocabulary-level prediction unit in the model’s output distribution, with each token associated with a logit indicating its selection likelihood.
-This approach inherently limits their ability to provide numerical probability scores, which are crucial for assessing model confidence.
-Moreover, general-purpose LMMs are not explicitly trained as classifiers with numerical labels, making direct probability estimation an unpreferred methodology.
-To address this limitation, FakeScope is trained using binary authenticity judgments and incorporates a token soft-scoring strategy to convert model outputs into quantitative probability estimates.
-In specific, as depicted in Fig.
-5, we define fake and real as anchor tokens and apply softmax normalization over their corresponding logits within a standardized prompting framework:
-USER: <Image> What is the authenticity of this image?
-ASSISTANT: This image is [auth_token].
-To compute probabilities, we extract the logits at the specific position of [auth_token], denoted by $\mathcal{T}=\{\tilde{t}_{j}|_{j=0}^{1}\}=\{\tilde{t}_{fake},\ \tilde{t}_{real}\}$, where $\tilde{t}_{(\cdot)}$ represents case variations of the anchor tokens to enhance robustness.
-The probability $p_{i}$ of each class is then calculated as follows,
-$p_{j}=\frac{e^{\sum_{v\in\tilde{t}_{j}}z^{(v)}}}{\sum_{j=0}^{1}e^{\sum_{v\in\tilde{t}_{j}}z^{(v)}}},$ (2)
-where $z^{(v)}$ represents the logit associated with token $v$.
-By leveraging token-level probability distributions rather than solely selecting the most probable token, this approach retains nuanced confidence signals that would otherwise be discarded during deterministic text generation.
-This enables FakeScope to generate quantifiable probability estimates, improving its interpretability as a forensic detector.
 VI Experiments and Analyses
@@
   • Generalization Capability: We examine how well FakeScope generalizes to unseen image content and generators (Sec. VI-D).
-Additionally, we perform ablation studies to analyze the impact of different training data components and scales (Sec. VI-E). Detailed experimental setups are provided below.
+Additionally, we perform ablation studies to analyze the impact of different training data components, scales, and inference strategies (Sec. VI-E). Detailed experimental setups are provided below.
 VI-A1 Implementation Details
-Training FakeScope is implemented with PyTorch on eight Nvidia H20-NVLink 98G GPUs, and the evaluations are conducted on one NVIDIA Tesla A100 40G GPU.
-All training are set to full-schedule mode, with detailed parameter configurations provided in SM.
-To ensure FakeScope serves as a unified forensic model for AI-generated images, we retain a single checkpoint per base model for all evaluations, ensuring fair comparisons.
-Furthermore, to prevent data leakage, no images from the testing datasets are reused in the training data.
-To ensure fairness, no data augmentation strategy is employed throughout the entire training process of FakeScope.
+FakeScope is trained in PyTorch on eight NVIDIA H20-NVLink 98GB GPUs, and evaluated on a single NVIDIA Tesla A100 40GB GPU.
+All runs use the full-schedule setting with full-parameter fine-tuning of the language model and modality projector, while the vision encoder is frozen for efficiency and stability.
+Detailed hyperparameters are provided in the SM.
+Training is performed for one epoch and completes in approximately 34 hours.
+For each base model, we retain a single checkpoint for all evaluations to ensure fair comparisons.
+No data augmentation is applied during training.
 VI-A2 Evaluation Datasets
-We evaluate the proposed FakeScope’s multifaceted performance on five commonly used datasets in the area of AI-generated image detection.
-As listed in Table II, we validate multimodal forensic capabilities on FakeBench [7], covering both binary authenticity and multimodal forensic criteria.
-To assess generalization to unseen content and generation models, we use AntifakePrompt [71], which contains fake images derived from real Microsoft COCO [74] images that are unseen during training.
-Additionally, we employ the non-overlapping subset of AIGCDetectBenchmark [72], excluding samples overlapping with ImageNet [62] to avoid contamination and ensure fair generalization evaluation.
-We further test FakeScope on user-generated images from social networks to support in-the-wild evaluation, using the WildRF [73] and SynthWildX [4] datasets, where authenticity labels are derived from social tags.
-It is worth noting that the scenarios like deepfakes (e.g.
-the FF++ dataset [75]) and are treated as out-of-distribution (OOD) cases, given their fundamental difference from fully AI-generated images.
-Nonetheless, they are included in our generalization evaluation to assess the model’s adaptability.
+We evaluate FakeScope on five widely used benchmarks for AI-generated image detection.
+As summarized in Table I, we assess multimodal forensic capabilities on FakeBench [40], which covers both binary authenticity prediction and open-ended forensic criteria.
+We further extend FakeBench to FakeBenchMCQ with 3,000 close-ended multiple-choice questions on local visual attributes to evaluate fine-grained forensic attribution; details are provided in the SM.
+To assess generalization to unseen content and generation models, we use AntifakePrompt [8], whose fake images are derived from Microsoft COCO [42] exclusive in training data.
+Additionally, for AIGCDetectBenchmark [82], we employ the strictly non-overlapping evaluation split by removing all real images from ImageNet [17] and fake images whose underlying real-image provenance traces back to ImageNet.
+This source-level isolation ensures no content is shared across training and testing.
+For in-the-wild evaluation, we use WildRF [7] and SynthWildX [12], which consist of user-generated social-media images.
+Finally, scenarios such as deepfakes (e.g., FF++ [62]) and image super-resolution are treated as out-of-distribution (OOD) cases because they typically modify real images rather than performing end-to-end synthesis; we nevertheless include them in the generalization evaluation.
 VI-A3 Baseline Models
-The proposed FakeScope is compared with a diverse set of models, including open-sourced LMMs (InstructBLIP [76], IDEFICS-Instruct [77], InternLM-XComposer2-vl [31], LLaVA-v1.5 [36], Qwen-VL [78], mPLUG-Owl2 [37]), proprietary LMMs (GPT-4V [33], GeminiPro [34], Claude3-Sonnet [35]), data-driven binary forensic models (CNNSpot [20], UnivFD [3], FreDect [22]), and the vision-language contrastive model CLIP-ViT-Large-14 [79].
-For detection evaluation, LMMs supports both qualitative (textual judgments) and quantitative (probabilistic estimation) settings, whereas binary detectors are limited to the quantitative mode.
-To ensure fair comparisons, the data-driven models are retrained on the same images as FakeScope without data augmentation, and all tested LMMs are evaluated using same prompting templates as FakeScope.
-These settings remain consistent across all experiments unless explicitly stated otherwise.
-TABLE III: Forensic detection results on the FakeBench-FakeClass [7] dataset under the qualitative setting.
-The detection Acc (%) of each model is reported.
+FakeScope is compared against a broad set of baselines: six open-source LMMs, including InstructBLIP [14], IDEFICS-Instruct [39], InternLM-XComposer2-VL [20], LLaVA-v1.5 [45], Qwen-VL [4], and mPLUG-Owl2 [78]; three proprietary LMMs, namely GPT-4V [54], GeminiPro [67], and Claude3-Sonnet [3]; three data-driven binary detectors, CNNSpot [68], UnivFD [53], and FreDect [24]; and the vision-language contrastive model CLIP-ViT-Large-14 [57].
+For detection, LMMs are evaluated in both qualitative (textual judgment) and quantitative (probability estimation) settings, while binary detectors are evaluated only in the quantitative setting.
+To ensure fairness, the data-driven detectors are retrained on the same images as FakeScope.
+All LMMs are evaluated using the same prompting templates as FakeScope.
+These settings remain consistent across all the experiments, unless explicitly stated otherwise.
+TABLE II: Forensic detection results on the FakeBench-FakeClass [40] under the qualitative setting.
+The detection Acc.
+(%) of each model is reported.
 The first and second best LMMs are highlighted in bold and underlined, respectively.
@@
 VI-A4 Performance Measures
-FakeScope is versatile in coping with forensic detection and in-depth analysis simultaneously.
-To comprehensively evaluate its multidimensional capabilities, we adopt a range of metrics.
+FakeScope is designed to handle both forensic detection and open-ended analysis, so we evaluate it with a diverse set of metrics.
 For forensic detection, we consider both qualitative and quantitative settings.
-Specifically, qualitative classification accuracy is measured by $Acc=N_{c}/N_{t}$, where $N_{c}$ and $N_{t}$ denote the number of correct and total predictions, respectively.
-For quantitative assessment, we follow [20, 3] and use threshold-independent calibration metrics, including average precision (AP) and area under the receiver operating characteristic curve (AUC), which are robust to class imbalance across datasets.
-To assess transparency-related performance, we evaluate the alignment between model responses and reference answers by employing both conventional word-level metrics (i.e., BLEU [80], ROUGE [81], and BERTScore [82]) and the LLM-as-a-judge strategy [83] to measure semantic-level alignment, following [40, 36, 7, 16].
-The measures are then normalized and averaged to obtain the overall reflection, where higher values indicate better performance on transparency-related abilities.
-Details on evaluation measures are provided in SM.
+In the qualitative setting, classification accuracy is defined as $Acc.=N_{c}/N_{t}$, where $N_{c}$ and $N_{t}$ are the numbers of correct and total responses.
+In the quantitative setting, following [68, 53], we use threshold-independent calibration metrics, including average precision (AP) and area under the ROC curve (AUC), which are more robust to dataset class imbalance.
+To assess transparency-related performance, we measure how well model responses align with reference answers both conventional word-level metrics (BLEU [55], ROUGE [10], and BERTScore [59]) and the LLM-as-a-judge strategy [26] to measure semantic-level alignment, following [50, 45, 40, 71].
+These scores are normalized and averaged into an overall transparency score, where higher values indicate better performance.
+Forensic attribution is evaluated by accuracy on FakeBenchMCQ.
+Details on the evaluation measures are further provided in the SM.
+VI-B Multimodal Forensic Capability
 Figure 6: Forensic detection results on FakeClass [40] under the quantitative setting.
-Calibration measures Acc (threshold is set to 0.5 for balanced indication), AP, and AUC (%) are reported for each model.
+AP and AUC (%) are reported for each image generation model.
 The FakeScope far exceeds the other models on each subset and the entire dataset.
-VI-B Multimodal Forensic Capability
-The multimodal forensic capabilities of the proposed FakeScope model are quantitatively assessed on FakeBench containing FakeClass, FakeClue, and FakeQA subsets with respect to four ability criteria for LMMs comprehensively defined in [7], which are stated as follows.
+The multimodal forensic capabilities of the proposed FakeScope model are quantitatively assessed on FakeBench containing FakeClass, FakeClue, and FakeQA subsets concerning four ability criteria for LMMs comprehensively defined in [40], which are stated as follows.
   • Detecting, which evaluates whether forensic classifiers can accurately differentiate AI-generated fake images from real ones;
@@
   • Reasoning, which assesses whether LMMs can identify generative visual trace evidence and correctly infer the unknown image authenticity;
-  • Fine-grained Analyzing, which investigates whether LMMs can discuss and analyze specific aspects of telltale visual clues.
+  • Fine-grained Analyzing, which investigates whether LMMs can discuss and analyze specific aspects of telltale visual clues;
+  • Forensic Attributing, which investigate whether LMMs can spot specific forensic visual cues in images.
 In practice, all LMM responses are generated using the greedy search strategy to ensure reproducibility.
 The quantitative evaluation results for the four criteria are presented below, with qualitative examples provided in SM.
-TABLE IV: Multimodal forensic capability comparisons on FakeBench [7], including detecting on FakeClass, interpreting on FakeClue-faultfinding, reasoning on FakeClue-inference, and fine-grained forgery analyzing on FakeQA.
+TABLE III: Multimodal forensic capability comparisons on FakeBench [40], including detecting on FakeClass, interpreting on FakeClue-faultfinding, reasoning on FakeClue-inference, and fine-grained forgery analyzing on FakeQA.
+The -M and -L denote base model mPLUG-Owl2 and LLaVA-1.5, respectively.
 The B.-1/B.-2 denote BLEU-1/2, R.-L is ROUGE-L, Sim.
 is BERTScore similarity, and $\alpha$, $\rho$, $\kappa$ correspond to completeness, preciseness, and relevance judged by LLMs.
-The first- and second-best performances of each column are highlighted in bold and underlined, respectively.
+Each column’s first- and second-best performances are highlighted in bold and underlined.
 Capabilities Interpreting Reasoning Fine-grained Analyzing Overall
-LMMs Word-level Semantic-level Avr.$\uparrow$ Word-level Semantic-level Avr.$\uparrow$ Word-level Semantic-level Avr.$\uparrow$
+[…]
 B.-1 B.-2 R.-L Sim. $\alpha$ $\rho$ $\kappa$ B.-1 B.-2 R.-L Sim. $\alpha$ $\rho$ $\kappa$ B.-1 B.-2 R.-L Sim. $\alpha$ $\rho$ $\kappa$
@@
 Improvement +0.289 +0.207 +0.211 +0.242 +0.081 +0.689 +0.266 +0.172 +0.286 +0.211 +0.214 +0.280 +0.968 +0.838 +0.610 +0.299 +0.176 +0.133 +0.144 +0.251 +0.388 +0.645 +0.220 +0.159 +0.210
-1) Detection Task: Simple detection of fake and real images is the fundamental capability of forensic detectors.
-In Table III, we compare the performance of the proposed FakeScope with other LMMs under the qualitative setting on the FakeClass dataset.
-The FakeClass dataset contains 6,000 diversified query-answer pairs for judging absolute image authenticity with balanced authenticity proportions and adequate query diversity.
-The detection accuracy is measured by the correctness of responses to these queries.
-As reflected in the results, FakeScope of the two versions simultaneously exhibits noticeable improvements in detection reliability across all generators, authenticity categories, and query types, suggesting that FakeInstruct can significantly enhance the forensic awareness of LMMs on binary classification.
-Notably, FakeScope achieves a peak accuracy of 99.68%, outperforming even the best-performing human expert, and marking a significant milestone in automated forensic analysis.
-In Fig.
-6, we further present the performance comparison using calibration measures under the quantitative setting.
-In this case, one can still witness a consistently significant advantage across all categories over both other LMMs, indicating the effectiveness of the proposed probability estimation strategy based on token soft-scoring.
-Notably, since we do not explicitly incorporate numerical values as the supervision when training LMMs as forensic detectors, the results further suggest that LMMs can effectively adapt to hard authenticity labels via the internalized probabilistic reasoning process and perform well on estimating numerical probabilities.
-In particular, despite being trained with the same image data, CNNSpot [20], UnivFD [3], and FreDect [22] exhibit inferior performance compared with FakeScope.
-While these data-driven methods can be effective in some scenarios, their performance often benefits from elaborate data augmentation [20].
-In contrast, under the controlled setting of this study where no image-level augmentation is applied, purely training with binary labels may fall short in capturing the intricacies of forensic cues.
-This further underscores the effectiveness of the multimodal information in FakeInstruct and highlights LMMs’ superiority to integrate multimodal information for AI-generated image forensics.
-2) Interpreting Task: We find out that incorporating visual trace evidence into training can significantly enhance the interpretation ability of LMMs.
-In practice, LMMs are evaluated using the fault-finding prompting mode of FakeClue, where models are instructed to explain image authenticity based on visual clues, such as in the prompt “This is a fake image generated by AI, explain the reasons”.
-In the first category of Table IV, the two variants of FakeScope demonstrate substantial improvements in interpreting ability, especially the word-level similarities and semantic relevance level to the ground truth.
-These findings indicate that training on FakeInstruct conspicuously enhances the effect-to-cause awareness concerning the forensic attributes of LMMs.
-3) Reasoning Task: Forensic reasoning is the inverse of interpretation, assessing LMMs’ inference capability in deriving authenticity judgments from visual trace evidence.
+TABLE IV: Forensic attribution capability (%) on FakeBenchMCQ. The -M and -L denote base model mPLUG-Owl2 and LLaVA-1.5.
+Model Question Type Attribution Level Overall
+Yes/No What Where How Low Medium High
+Random guess 50.13 37.60 24.27 26.93 34.50 37.80 31.90 34.73
+GPT-4V (Teacher) 79.47 68.80 55.87 62.40 70.70 61.50 67.60 66.60
+GeminiPro 70.37 66.20 52.13 70.80 65.20 62.60 66.80 64.88
+Claude3 Sonnet 68.27 67.73 66.80 65.20 68.30 68.80 63.90 67.00
+InstructBLIP 56.27 56.00 54.93 53.60 57.70 58.40 49.50 55.20
+IDEFICS-Instruct 48.67 49.60 52.27 53.47 50.20 49.70 53.10 51.00
+InternLM-XC.2-vl 69.73 68.13 66.27 59.47 66.30 65.50 65.90 65.90
+LLaVA-v1.5 60.40 62.27 63.07 61.60 63.80 62.60 59.11 61.83
+Qwen-VL 64.13 64.93 60.93 60.67 62.30 61.80 63.90 62.67
+mPLUG-Owl2 70.53 66.93 62.27 64.13 64.80 66.30 66.80 65.97
+FakeScope-M 75.47 79.73 78.53 80.27 76.80 80.50 78.20 78.50
+Improvement +4.94 +12.80 +16.26 +16.14 +12.00 +14.20 +11.40 +12.53
+FakeScope-L 79.47 78.00 78.27 75.20 82.80 73.70 76.70 77.74
+Improvement +19.07 +15.73 +15.20 +13.60 +19.00 +11.10 +17.59 +15.91
+1) Detection Task: Distinguishing real from fake images is the core capability of forensic detectors.
+We compare FakeScope with other LMMs under the qualitative setting using the FakeClass dataset, which includes 6,000 diverse QA pairs with balanced authenticity labels and varied queries.
+As shown in Table II, both versions of FakeScope consistently outperform other models across all subcategories, indicating that FakeInstruct notably enhances LMMs’ forensic detection capabilities.
+FakeScope reaches a peak accuracy of 99.68%, exceeding even the strongest human expert.
+Figure 6 further compares models under the quantitative setting using calibration-based metrics.
+FakeScope remains the strongest model, demonstrating the effectiveness of the token-based soft-scoring strategy.
+Notably, despite receiving no explicit numerical supervision during training, it aligns closely with hard labels and produces reliable probability estimates, suggesting strong internalized reasoning.
+Even when trained on the same image data, conventional detectors perform worse than FakeScope.
+Without image-level augmentation, these data-driven detectors degrade markedly [68], underscoring the value of FakeInstruct’s multimodal supervision and the stronger forensic integration ability of LMMs.
+2) Interpreting Task: We evaluate LMMs with the fault-finding prompting mode of FakeClue, where models are instructed to explain image authenticity based on visual descriptions, such as in the prompt “This is a fake image generated by AI, explain the reasons”.
+In the first category of Table III, FakeScope achieve clear gains in this task, which indicates that training on FakeInstruct conspicuously enhances the effect-to-cause awareness concerning the forensic attributes of LMMs.
+3) Reasoning Task: The forensic reasoning task is the inverse of interpretation, assessing LMMs’ inference capability in deriving authenticity judgments from visual trace evidence.
 This task is evaluated using the prompts in the inference mode of FakeClue, such as “Inspect this image carefully and formulate an evaluation regarding its authenticity”.
-As observed in the experimental results listed in the second category of Table IV, while all baseline LMMs generally exhibit weak forensic reasoning capabilities, the proposed FakeScope achieves a substantial improvement after training on FakeInstruct and even surpasses all compared models, reaching SOTA performance.
-Most notably, its reasoning capability also outperforms the teacher model GPT-4V [33] by over 30%, demonstrating the effectiveness of FakeInstruct in enhancing cause-to-effect forensic awareness and triggering forensic reasoning in LMMs.
-4) Fine-Grained Analyzing Task: To further evaluate the effectiveness of FakeScope in analyzing specific aspects of image forgery, we conduct experiments on the FakeQA dataset, which consists of queries regarding fine-grained visual forgeries.
-As shown in the third category of Table IV, FakeScope exhibits notable improvements over both baseline models and its teacher model, highlighting its enhanced fine-grained forensic analyzing capabilities.
-However, despite these promising results, this ability still shows room for improvement compared to the other two transparency-related forensic capabilities interpreting and reasoning.
-This suggests potential avenues for further optimization, particularly in refining its sensitivity to subtle forgery artifacts.
-TABLE V: In-the-wild forensic detection ability on WildRF [73] and SynthWildX [4] containing user-generated image content and unknown generators.
-The AP and AUC (%) are reported, and the first- and second-best performers for each category are highlighted in bold and underline respectively.
+As shown in the second category of Table III, baselines generally perform poorly on this task, whereas FakeScope improves substantially after training on FakeInstruct and achieves state-of-the-art results.
+Most notably, it surpasses the teacher model GPT-4V by more than 30%, highlighting FakeInstruct’s effectiveness in strengthening cause-to-effect forensic awareness.
+4) Fine-Grained Analyzing Task: We assess FakeScope on FakeQA, which targets fine-grained visual forgery queries.
+As shown in the third category of Table III, FakeScope clearly outperforms both baseline and teacher models, demonstrating stronger fine-grained forensic analysis.
+However, this capability still lags behind its interpreting and reasoning performance, suggesting the future potential to improve its answering open-ended questions on subtle forgery artifacts.
+Figure 7: Accuracy as a function of detection threshold.
+For each model, we present the relations between the detection threshold (x-axis) and accuracy (y-axis) on all generators of the five testing databases.
+The FakeScope of two versions of base models (the last two sub-figures) is the only one that can achieve a great bargain between threshold sensitivity and detection accuracy, indicating remarkable robustness.
+TABLE V: In-the-wild performance comparison. The AP and AUC (%) are reported, and top two performers for each category are highlighted in bold and underline respectively.
 Datasets WildRF SynthWildX Average
@@
 FakeScope (Ours, LLaVA-v1.5) 80.01/ 78.60 67.94/ 73.23 75.48/ 76.09 87.37/ 87.44 78.94/ 81.76 81.99/ 83.40 78.46/ 80.09
-TABLE VI: Generalization capability validated on AntifakePrompt [71] and AIGCDetectBenchmark [72].
-The AP (%) of various classifiers across 17 individual generators are reported.
-Chance is 50% and best possible performance is 100%.
-All the LMMs are tested under the quantitative setting.
+TABLE VI: Generalization capability comparison.
+The AP (%) of various classifiers across 17 individual generators is reported.
 Values in gray represent the generators seen in training.
-No content overlapping exists in the results.
-The best model in each column is highlighted in bold.
+The top performer is highlighted in bold.
 Datasets AIGCDetectBenchmark AntifakePrompt Unseen Average
 […]
-CNNSpot 64.86 61.41 70.33 60.10 63.08 56.41 70.19 82.85 92.51 86.50 56.61 60.25 48.88 52.15 50.54 50.22 95.15 60.71 66.00
-UnivFD 61.37 50.36 39.78 55.61 48.49 51.91 56.22 43.25 35.75 44.24 45.36 58.17 50.85 57.46 80.99 60.68 68.42 57.24 53.47
-FreDect 49.93 55.13 49.99 53.26 54.93 49.97 53.61 29.79 37.78 38.36 35.80 44.46 49.86 49.93 48.02 49.81 56.10 50.63 47.45
-CLIP-ViT-Large-14 89.22 71.12 94.92 88.58 64.40 81.17 60.43 29.20 31.87 37.28 49.44 70.85 62.86 63.26 85.15 71.71 41.85 71.38 64.31
-InternLM-XC.2-vl 52.15 46.40 61.16 50.47 52.28 44.29 68.72 35.73 41.03 48.59 45.49 62.15 52.39 58.33 60.58 58.46 71.70 58.23 53.52
-InstructBLIP 50.27 40.79 32.31 88.63 40.42 49.36 55.46 25.98 29.87 25.86 72.11 64.04 59.18 74.46 95.64 81.84 97.48 67.17 57.86
-Qwen-VL 72.27 42.91 41.39 82.54 44.80 48.92 84.18 75.39 69.55 72.97 83.83 88.15 66.50 78.31 93.55 89.05 91.31 73.52 72.10
-IDEFICS-Instruct 40.26 51.87 53.68 39.40 47.26 66.99 35.16 22.03 23.55 25.85 22.45 41.24 46.53 41.11 34.81 44.27 48.49 45.36 40.29
-mPLUG-Owl2 58.44 48.91 83.24 51.39 45.26 41.84 69.79 45.10 46.10 44.71 55.03 74.34 50.71 57.82 77.64 68.28 71.16 62.86 58.22
-LLaVA-v1.5 40.97 44.09 41.30 38.08 43.03 56.01 42.36 24.32 25.07 26.71 32.78 45.13 47.93 45.08 37.69 43.91 80.87 47.40 42.08
+CNNSpot [68] 64.86 61.41 70.33 60.10 63.08 56.41 70.19 82.85 92.51 86.50 56.61 60.25 48.88 52.15 50.54 50.22 95.15 60.71 66.00
+UnivFD [53] 61.37 50.36 39.78 55.61 48.49 51.91 56.22 43.25 35.75 44.24 45.36 58.17 50.85 57.46 80.99 60.68 68.42 57.24 53.47
+FreDect [24] 49.93 55.13 49.99 53.26 54.93 49.97 53.61 29.79 37.78 38.36 35.80 44.46 49.86 49.93 48.02 49.81 56.10 50.63 47.45
+CLIP-ViT-Large-14 [57] 89.22 71.12 94.92 88.58 64.40 81.17 60.43 29.20 31.87 37.28 49.44 70.85 62.86 63.26 85.15 71.71 41.85 71.38 64.31
+InternLM-XC.2-vl [20] 52.15 46.40 61.16 50.47 52.28 44.29 68.72 35.73 41.03 48.59 45.49 62.15 52.39 58.33 60.58 58.46 71.70 58.23 53.52
+InstructBLIP [14] 50.27 40.79 32.31 88.63 40.42 49.36 55.46 25.98 29.87 25.86 72.11 64.04 59.18 74.46 95.64 81.84 97.48 67.17 57.86
+Qwen-VL [4] 72.27 42.91 41.39 82.54 44.80 48.92 84.18 75.39 69.55 72.97 83.83 88.15 66.50 78.31 93.55 89.05 91.31 73.52 72.10
+IDEFICS-Instruct [39] 40.26 51.87 53.68 39.40 47.26 66.99 35.16 22.03 23.55 25.85 22.45 41.24 46.53 41.11 34.81 44.27 48.49 45.36 40.29
+mPLUG-Owl2 [78] 58.44 48.91 83.24 51.39 45.26 41.84 69.79 45.10 46.10 44.71 55.03 74.34 50.71 57.82 77.64 68.28 71.16 62.86 58.22
+LLaVA-v1.5 [45] 40.97 44.09 41.30 38.08 43.03 56.01 42.36 24.32 25.07 26.71 32.78 45.13 47.93 45.08 37.69 43.91 80.87 47.40 42.08
 FakeScope (mPLUG-Owl2) 89.70 92.96 88.63 89.83 86.94 93.32 98.20 97.27 99.17 98.80 99.04 99.77 98.20 98.99 99.72 98.80 99.96 95.67 95.84
 FakeScope (LLaVA-v1.5) 99.92 98.01 100.00 99.66 98.01 99.45 98.96 98.21 99.04 99.10 99.04 99.65 82.59 95.97 96.40 97.20 73.64 94.68 96.17
+5) Forensic Attribution Task: We evaluate forensic attribution on FakeBenchMCQ, containing close-ended questions on specific visual evidence.
+As listed in Table IV, FakeScope achieves clear improvements over its base models, indicating that it can not only justify detection with long-form rationales but also identify concrete forensic evidence effectively.
+This further confirms that FakeInstruct successfully imparts visual forensic awareness to LMMs.
+Among all dimensions, the largest gain appears on where questions.
+We also observe larger gains on low- and medium-level attributes, suggesting the model could potentially be further improved by incorporating more conceptually high-level forensic knowledge.
 VI-C In-the-wild Performance
-To evaluate the real-world applicability of our model, we conduct in-the-wild experiments on diverse, unstructured image data collected from social media, including WildRF [73] and SynthWildX [4].
-These datasets contains real-world samples with various distortions, such as compression artifacts, post-processing modifications, and noise, which pose additional challenges for forensic models.
-The testing protocol aligns with other experiments, and all LMMs are evaluated in the quantitative setting.
-The AP and AUC values for detection are listed in Table V.
-As reflected by the results, FakeScope equipped with mPLUG-Owl2 [37]consistently outperforms other state-of-the-art methods across the majority of generators, achieving superior overall performance in the in-the-wild testing.
-Notably, it demonstrates competitive results on images generated by Dalle3 and Midjourney, ranking second only to CNNSpot [20].
-These findings highlight FakeScope’s robustness and effectiveness in handling unconstrained image content and generation models, reinforcing its adaptability to real-world forensic challenges beyond in-lab benchmarks.
-Figure 7: Accuracy as a function of detection threshold.
-For each model, we present the relations between the detection threshold (x-axis) and accuracy (y-axis) on all generators of five testing databases.
-It turns out the FakeScope of two versions of base models (the last two sub-figures) is the only one that can achieve a great bargain between threshold sensitivity and detection accuracy, indicating remarkable robustness.
-Figure 8: Performance on FakeBench [7] with respect to training data scale, demonstrating the effectiveness of FakeInstruct data.
-The x-axis indicates the proportion of FakeInstruct used for training and the y-axis is the corresponding performance indicator of each ability dimension.
+To assess real-world applicability, we evaluate on in-the-wild datasets built from social-media imagery, including WildRF [7] and SynthWildX [12].
+These contain real-world degradations which pose additional challenges for forensic models.
+As reflected by the results in Table V, FakeScope equipped with mPLUG-Owl2 [78] consistently outperforms prior state-of-the-art methods across most generators and achieves the best overall in-the-wild performance.
+It is also competitive on Dalle3 and Midjourney, ranking second only to CNNSpot [68].
+These findings highlight FakeScope’s robustness in handling real-world forensic challenges beyond in-lab benchmarks.
 VI-D Generalization Capability
-For forensic detectors, the generalization ability on unacquainted image contents and generators is crucial for practical deployment.
-To assess this, we conduct a more challenging experiment to evaluate whether FakeScope is well generalizable on unseen image content and generators.
-Practically, we test all methods on the uncontaminated content proportions of AntifakePrompt [71] and AIGCDetectBenchmark [72], in comparison with the data employed in FakeInstruct.
-In particular, the AntifakePrompt dataset covers several OOD tasks beyond the primary scope of this work, such as image super-resolution, face spoofing [75], and inpainting, which pose a greater challenge to model generalization and robustness.
-In Table VI, the AP and AUC scores of each model across 17 image generators are presented, including 6 seen and 11 unseen ones.
-We also present the average AP over unseen generators as well as over all generators.
-From Table VI, it can be observed that the two variants of the FakeScope can achieve SOTA performance across all 17 generators and, notably, outperforms all baselines on the set of unseen generators.
-These results demonstrate that FakeScope can effectively handle novel image content and generation techniques, indicating strong generalization capability.
-Furthermore, we analyze model stability by plotting detection accuracy as a function of the decision threshold ($\in(0,1)$) for each method, thereby assessing their sensitivity to threshold selection across different generators.
+The generalization to unacquainted image contents and generators is crucial for practical deployment.
+To assess this, we test all methods on the uncontaminated content proportions of AntifakePrompt [8] and AIGCDetectBenchmark [82], whose content is disjoint from the data used in FakeInstruct.
+AntifakePrompt further includes OOD tasks beyond our primary scope, such as image super-resolution, face spoofing [62], and inpainting, which stress both robustness and generalization.
+Table VI presents results across 17 image generators, including six unseen and 11 unseen ones.
+Both FakeScope variants achieve superiority across all 17 generators and outperform all baselines on the unseen subset, indicating strong generalization to novel content and generation techniques.
+We further evaluate stability by plotting detection accuracy as a function of the decision threshold ($\in(0,1)$), measuring sensitivity to threshold selection across generators.
 In Fig.
-7, a separate curve is shown for each forensic detector by averaging detecting accuracy across all generators in each dataset.
-Unlike AUC and mAP, the accuracy (Acc) is highly sensitive to threshold variation, and models with unstable performance may suffer under suboptimal thresholds, leading to less robustness.
-It can be observed from Fig.
-7 that both versions of FakeScope exhibit highly consistent peak accuracy across different datasets and maintain stable performance over a broad range of threshold values.
-In contrast, other models show considerable variability in optimal accuracy across datasets and are overly sensitive to specific threshold settings.
-This indicates that FakeScope possesses prominent robustness towards forensic threshold across different datasets and generators, thereby enhancing its applicability in real-world scenarios.
+7, each curve is obtained by averaging accuracy over all generators in each dataset.
+Unlike AUC and mAP, accuracy is highly threshold-dependent; unstable detectors can degrade substantially under suboptimal thresholds.
+As shown in Fig.
+7 both FakeScope variants exhibit consistent peak accuracy and remain stable over a wide threshold range.
+In contrast, competing methods show larger shifts in their optimal thresholds and greater performance volatility.
+This suggests FakeScope is less sensitive to threshold choice across datasets and generators, improving robustness in real-world settings.
 VI-E Ablation Studies
-To investigate the influence of different data components and scales on training FakeScope, we further conduct ablation studies on the influence of three aspects during tuning with mPLUG-Owl2 [37]: (1) scale of training data; (2) collaboration of FakeChain and visual instructions; and (3) participation of multimodal contextual information.
-1) Effects of Training Data Scale: To effectively handle data-scarce scenarios is crucial for forensic models, yet it remains increasingly challenging due to the high cost and difficulty of acquiring high-quality training data.
-To evaluate the effectiveness of FakeInstruct dataset, we randomly sample four subsets—80%, 60%, 40%, and 20%—while preserving the original data composition, and analyze the impact on training outcomes.
-As illustrated in Fig.
-8, the resulting FakeScope variants exhibit remarkable resilience to performance degradation as the available training data decreases, demonstrating the strong utility of FakeInstruct.
-Notably, the transparency-related performance remains a significant improvement over the baselines, even at the extreme 20% data proportion, and continues to strengthen as the training data size grows.
-This trend highlights the effectiveness of the multimodal forensic knowledge embedded in FakeInstruct, further demonstrating that the reasoning data can substantially enhance the forensic capabilities of large multimodal models.
-2) Effects of Collaborative Training: During training, we combine different subsets to jointly optimize a unified forensic foundation model.
-To validate the effectiveness of this strategy, we compare it against task-specific training focused solely on transparent forensic capabilities.
-As indicated by the results listed in Table VII, the collaborative training scheme—integrating long-form reasoning with granular visual instructions—yields significant improvements in model performance.
-Although the various subsets in FakeInstruct originate from a shared data source, this strategy remains cost-efficient and enables FakeScope to generalize more effectively across diverse forensic tasks, demonstrating greater stability across a broader knowledge domain.
+We analyze how data composition and scale affect FakeScope training with mPLUG-Owl2 [78].
+Specifically, we study: (1) instruction scale, i.e., the effect of dataset size; (2) instructional diversity, by comparing FakeChain-only training against enriched multimodal supervision from FakeInstruct; (3) multimodal awareness, via detection performance with and without multimodal context.
+Besides, we further validate the (4) effectiveness of the proposed token soft-scoring strategy.
+1) Effects of Training Data Scale: Data scarcity is a core bottleneck for forensic modeling because high-quality annotations are expensive to obtain.
+To assess the data efficiency of FakeInstruct, we sample 80%, 60%, 40%, and 20% subsets while preserving the original composition.
+As shown in Fig.
+8, FakeScope degrades gracefully as training data decreases, indicating strong robustness of FakeInstruct.
+Even at 20%, performance remains clearly above baselines, and improves steadily with larger subsets.
+This trend suggests that the gains are driven primarily by the forensic knowledge encoded in FakeInstruct rather than data volume alone.
+2) Effects of Instructional Diversity: We jointly train a unified forensic foundation model on multiple subsets and compare this setting with task-specific training focused only on transparent forensics.
+As shown in Table VII, collaborative training that combines long-chain reasoning with fine-grained visual instruction consistently improves performance, which benefits cross-task generalization.
+Full-schedule training improves all capabilities while reducing cross-task interference, indicating complementary rather than competitive interactions among forensic knowledge types.
+In particular, both analysis and attribution benefit from fine-grained supervision, whereas training with FakeChain alone provides no comparable improvement.
+These results validate the effectiveness of our visual instruction strategy.
+Figure 8: Performance on FakeBench [40] concerning training data scale.
+The x-axis indicates the proportion of FakeInstruct used for training, and y-axis shows corresponding performance indicator of each ability dimension.
 TABLE VII: Comparison on transparent forensic capabilities between only FakeChain and full FakeInstruct as training dataset. The best performance of each column is highlighted in bold.
-Training Strategy Detect Interpret Reason Analyze
-none (baseline) 58.22% 0.423 0.296 0.368
-only FakeChain 83.27% 0.593 0.576 0.350
-full FakeInstruct 99.68% 0.638 0.614 0.540
-TABLE VIII: Comparison on forensic detection ability between full FakeInstruct, only absolute judgment, and only contextual data.
-The AP (%) on each dataset is reported.
-The best performance of each column is highlighted in bold.
+Training Strategy Detect Interpret Reason Analyze Attribute
+none (baseline) 58.22% 0.423 0.296 0.368 65.97%
+only FakeChain 83.27% 0.593 0.576 0.350 68.43%
+full FakeInstruct 99.68% 0.638 0.614 0.540 78.50%
+TABLE VIII: Comparison on forensic detection ability among full FakeInstruct, only absolute judgment, and only contextual data containing trace evidence instructions.
+AP (%) on each dataset is reported.
 Dataset/ Training Strategy none (baseline) only absolute only contextual full (employed)
-FakeBench [7] 68.37 94.06${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+25.96}}}$ 84.13${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+15.76}}}$ 99.94${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+31.57}}}$
-AntiFakePrompt [71] 60.06 90.12${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+30.06}}}$ 74.84${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+14.78}}}$ 98.90${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+38.84}}}$
-AIGCDetecBenchmark [72] 54.84 81.65${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+26.81}}}$ 60.73${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+5.89}}}$ 90.23${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+35.39}}}$
-WildRF [73] 52.73 82.34${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+29.61}}}$ 62.54${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+9.81}}}$ 90.05${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+37.32}}}$
-SynthWildX [4] 47.41 82.78${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+35.37}}}$ 52.59${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+5.18}}}$ 86.93${}_{\text{{\color[rgb]{0.1171875,0.46484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.46484375,0.515625}+39.52}}}$
-3) Effects of Multimodal Awareness: A defining feature that sets FakeScope apart from conventional binary forensic detectors is its integration of rich multimodal contextual information on visual trace evidence during training.
-While these data are aimed at enhancing transparency-related forensic capabilities, they might affect the performance on simple detection.
-To validate whether such information also improves pure forensic detection, we compare the performance of models trained with and without it.
-As compared in Table VIII, incorporating contextual information not only improves the model’s transparency but also enhances detection accuracy and generalization, even when using the same amount of image data.
-In contrast, omitting absolute authenticity labels from training results in inconsistent improvements.
-These findings suggest that multimodal forensic cues can strengthen the model’s task awareness and fine-grained visual comprehension, enabling it to more effectively attend to forensic-relevant patterns.
-However, the model’s detection performance still heavily relies on absolute authenticity judgments, as relying solely on descriptive cues may introduce ambiguity into the decision-making process for authenticity.
+[…]
+[…]
+[…]
+[…]
+[…]
+3) Effects of Multimodal Awareness: A key distinction between FakeScope and conventional binary detectors is the use of multimodal evidence for visual trace analysis.
+Although this design is intended to improve transparent forensics, its contribution to basic detection is not obvious.
+We therefore compare models trained with and without trace-evidence instructions.
+As reported in Table VIII, adding multimodal context improves both detection accuracy and generalization, even under the same image budget.
+By contrast, removing absolute authenticity labels yields less stable gains.
+These results indicate that multimodal forensic cues enhance task awareness and fine-grained visual understanding, whereas robust detection still requires explicit authenticity supervision.
+TABLE IX: Ablations of the proposed token soft-scoring strategy: (i) auxiliary classifier; (ii) verbalized confidence; (iii) synonym ensembles for <auth_token>; and (iv) argmax vs.
+softmax scoring.
+We report mAP (%) on each dataset.
+Exp. Model Strategy Fake Bench Antifake Prompt AD Bench. WildRF Synth WildX
+(i) FakeScope (mPLUG-Owl2) coupled (ours) 99.94 98.90 90.23 90.05 86.93
+de-coupled 98.67 95.20 87.69 85.42 83.89
+FakeScope (LLaVA-v1.5) coupled (ours) 99.39 94.53 99.15 74.48 82.77
+de-coupled 96.78 95.99 96.45 60.80 79.76
+(ii) FakeScope (mPLUG-Owl2) token-based (ours) 99.94 98.90 90.23 90.05 86.93
+verbalizegreedy 66.08 60.43 50.87 52.36 45.42
+verbalizet=0.5 66.79 61.66 52.80 52.39 45.40
+verbalizet=1.0 65.32 58.75 51.43 50.28 43.27
+FakeScope (LLaVA-v1.5) token-based (ours) 99.39 94.53 99.15 74.48 82.77
+verbalizegreedy 68.89 62.58 52.66 54.90 49.87
+verbalizet=0.5 68.72 62.05 51.43 54.95 49.26
+verbalizet=1.0 66.77 61.02 51.35 52.73 46.58
+[…]
+[…]
+[…]
+[…]
+[…]
+fake+artificial+synthetic $\leftrightarrow$ real+genuine+authentic 97.78 97.29 90.30 89.03 86.81
+[…]
+[…]
+[…]
+[…]
+[…]
+fake+artificial+synthetic $\leftrightarrow$ real+genuine+authentic 99.43 92.88 99.02 72.65 83.21
+(iv) FakeScope (mPLUG-Owl2) softmax (ours) 99.94 98.90 90.23 90.05 86.93
+argmax 95.66 98.47 63.57 55.36 78.78
+FakeScope (LLaVA-v1.5) softmax (ours) 99.39 94.53 99.15 74.48 82.77
+argmax 93.55 95.64 85.55 52.92 77.42
+[…]
 VII Conclusions and Future Work
-In this paper, we propose a multimodal paradigm for transparent AI-generated image forensics, versatile in detection, in-depth analysis, and extended discussions.
-Based on the advanced cross-modal capability of large multimodal models (LMMs), we collect and leverage rich image forensic knowledge to enhance their forensic awareness.
-By emulating human causal inference based on visual trace evidence, we introduce a human-machine collaborative strategy and construct the first large-scale dataset containing structured forensic reasoning data.
-This reasoning foundation is further expanded into a million-scale visual instruction dataset, which is internalized into LMMs to improve their transparent forensic abilities on AI-generated images.
-Extensive experiments demonstrate consistent effectiveness of this paradigm across LMMs with varied meta frameworks.
-The proposed expert forensic model can reliably handle multiple forensic tasks under a unified paradigm, and exhibits strong generalization and in-the-wild capabilities.
-Besides, we discovered that LMMs can effectively align with hard authenticity labels and demonstrate impressive generalization on estimating numerical probabilities, even without explicit supervision.
-This work not only presents a viable pathway toward transparent AI-generated image forensics, but also offers insights into the development of trustworthy AI risk management within the broader multimodal landscape.
-Moreover, the proposed model could serve as the source of generating cross-modal forensic information, which may further support the development of more reliable forensic systems.
-Notwithstanding the use of advanced models such as GPT-4V to distill forensic knowledge, it may still represent an early-stage effort in the evolving trajectory of large multimodal models.
-While some uncertainties remain, this study nonetheless demonstrates the effectiveness of multimodal information and the feasibility of forensic transparency.
-We intend to further explore the abilities of newly emerged LMMs and contribute model update to community.
-Looking ahead, future efforts may focus on enhancing the model’s few-shot adaptability and dialectical reasoning capability for better robustness and explanatory power in more complex forensic contexts.
+This paper proposes a multimodal paradigm for transparent AI-generated image forensics that integrates detection, in-depth analysis, and extended discussions within a unified framework.
+To support this paradigm, we build FakeChain, a large-scale forensic reasoning dataset created through human-machine collaboration, and extend it into FakeInstruct, the first multimodal instruction-tuning dataset for image forensics with 2 million visual instructions.
+On this basis, we develop FakeScope, an expert large multimodal model that achieves superior performance in both closed-ended and open-ended forensic tasks, while showing strong generalization and robustness in the wild.
+Notably, FakeScope demonstrates zero-shot quantitative detection without explicit numerical supervision, suggesting an emergent capacity for probabilistic reasoning.
+More broadly, this work offers a practical path toward transparent and trustworthy AI-generated image forensics.
+Beyond accurate detection, FakeScope produces cross-modal forensic information, supporting explainable analysis.
+Our results also underscore the value of human-in-the-loop design and structured prompting for reliable knowledge distillation in LMMs.
+Although current progress still depends on advanced models such as GPT-4V and important uncertainties remain, the findings confirm the value of multimodal information for forensic transparency.
+Looking ahead, we intend to explore the abilities of newly emerged LMMs and contribute model updates to the community.
+Given the diminishing returns implied by scaling laws, future efforts may focus on enhancing few-shot adaptability and dialectical reasoning to boost robustness in more complex, unseen forensic scenarios.